### v1

# London WC2R 2LS, UK

Abstract
We demonstrate that all perturbative scale invariant heterotic sigma models with a compact target space $M^{D}$ are conformally invariant. The proof, presented in detail for up to and including two loops, utilises a geometric analogue of the $c$-theorem based on a generalisation of the Perelman’s results on geometric flows. Then, we present examples of scale invariant heterotic sigma models with target spaces that exhibit special geometry, which is characterised by the holonomy of the connection with torsion a 3-form, and explore the additional conditions that are necessary for such sigma models to be conformally invariant. For this, we find that the geometry of the target spaces is further restricted to be either conformally balanced or the a priori holonomy of the connection with torsion reduces. We identify the pattern of holonomy reduction in the cases that the holonomy is $SU(n)$ $(D=2n)$, $Sp(k)$ $D=4k)$, $G_{2}$ $(D=7)$ and $\mathrm{Spin}(7)$ $(D=8)$. We also investigate the properties of these geometries and present some examples.
1 Introduction
Conformal invariance in the context of 2-dimensional sigma models has been extensively explored because of its applications to string theory friedan ; CFMP ; Ts ; Ts2 ; CT . In particular, the conditions for a sigma model with metric $G$ and B-field couplings to be scale and conformal invariant have been described by Hull and Townsend in HullTownsend . More generally, the relationship between scale and conformal invariance in two-dimensional field theories has been explored by Polchinski Polchinski who established that all scale invariant unitary [^1] There are examples of non-unitary scale invariant theories that are not conformally invariant, see Polchinski2 page 260 and ISS . field theories with a discrete spectrum of operator scaling dimensions are conformally invariant. The proof is rather involved and uses the Zamolodchikov $c$-theorem Zam ; see DKST for a generalisation to four dimensions. Polchinski also noticed that scale invariant sigma models with only a metric coupling $G$ and compact target space $M^{D}$ are conformally invariant with constant dilaton. The proof relies on a global result established by Bourguignon Bourg in the context of geometric flows on manifolds.
The relation between geometric flows on manifolds and sigma model renormalisation group flows extends to the work of Perelman Perelman on the proof of the Poincaré conjecture who established an analogue of the $c$-theorem for geometric flows. The use of geometric flows to prove the Poincaré conjecture was initially proposed by Hamilton Hamilton , see MorganTian ; Hamilton2 ; Bamler ; Cao for reviews and further developments. The work of Perelman, adapted to sigma models, made the idea of interpreting the sigma model central charge as the spacetime effective action Tseytlin3 ; CP concise. These results were later extended and refined to theories with a $B$-field couplings by Oliynyk, Suneeta, and Woolgar OSW2 ; OSW , see also Tseytlin2 ; Huhu ; ST1 ; ST2 ; GFS for further work in both physics and geometry. In particular, OSW2 ; OSW ; Tseytlin2 proved the $c$-theorem for sigma models. This technology has been used in PapWitten to establish that scale invariant sigma models with both metric $G$ and $B$-field couplings and compact target space are conformally invariant. The latter proof relies on the invariance of the formalism under reparameterisations of the sigma model manifold $M^{D}$ and $B$-field gauge transformations. A related statement in the context of geometric flows has previously be given in GFS . By presenting counterexamples, in PapWitten it was also established that compactness, or at least geodesic completeness, of $M^{D}$ is a necessary and sufficient condition for scale invariant sigma models to be conformal.
Heterotic sigma models apart from the metric $G$ and $B$-field couplings, which are common with the models that have been examined so far, they also have a gauge field coupling $A$ GHMR . In addition, they are chiral and so they exhibit potential sigma model manifold reparameterisation and gauge transformation chiral anomalies. However, these are cancelled with assigning an anomalous variation to the $B$ field Hullwitten . Gravitational, gauge and mixed anomalies of certain supergravity theories also cancel with a similar mechanism previously proposed by Green and Schwarz mgjs . The understanding of these anomaly cancellations is essential to preserve these symmetries in the effective theory of heterotic strings and so ensure the geometric interpretation of the theory. In turn, the invariance of the heterotic string effective action under reparameterisations of $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations is key to the generalisation of the results of PapWitten to the heterotic sigma models.
The purpose of this article is to prove a geometric analogue of the $c$-theorem and use this to explore the relation between scale and conformal invariance in the context of heterotic sigma models. In particular, we shall establish that all perturbative scale invariant heterotic sigma models with compact target space are conformally invariant. This result has already been stated in PapWitten . Here, we shall supply the required formulae for the proof up to two loops in sigma model perturbation theory and explore the conditions on the geometry of sigma model target spaces required for conformal invariance. Then, we shall explore the relationship between scale and conformal invariance for a class of heterotic sigma models that have target spaces manifolds with special geometry – for an early related work see Hull . In particular, we shall demonstrate that heterotic sigma models with target spaces $D$-dimensional manifolds, $M^{D}$, for which the holonomy of the connection [^2] Note that ^$\widehat{\nabla}_{J}V^{I}\equiv D_{J}V^{I}+\frac{1}{2}H^{I}{}_{JK}V^{K}$, where $D$ is the Levi-Civita connection of $G$ and $H$ is given in (5), i.e. ${\mathrm{d}}H\not=0$. , $\widehat{\nabla}$, with skew-symmetric torsion $H$ is included in
$SU(n)~{}~{}(D=2n)\,,\,Sp(k)~{}~{}(D=4k)\,,\,G_{2}~{}~{}(D=7)\,,\,\mathrm{Spin}(8)~{}~{}(D=8)\,.$ (1)
and the curvature $F$ of $A$ satisfies an instanton-like condition are scale invariant up to two loops in perturbation theory. This choice of holonomy groups can be motivated from either spacetime or worldsheet considerations. In the former case, they arise as conditions on the geometry of $M^{D}$ for the existence of non-trivial solutions of the gravitino Killing spinor equation. The same applies for the instanton-like condition on $F$ – it arises as a condition for solving the gaugino Killing spinor equation, for a review on the geometry of supersymmetric heterotic backgrounds see GGP . In the latter case, the $\widehat{\nabla}$-covariantly constant forms [^3] Each of the holonomy groups in (1) is characterised by the existence of $\widehat{\nabla}$-covariantly constant forms on $M^{D}$, the fundamental forms of the holonomy group. on $M^{D}$ that characterise these holonomy groups generate transformations HowePap that leave invariant the heterotic sigma model action (2). The instanton-like condition on $F$ is required for the invariance of the gauge sector deLaOssa .
On comparing the conditions for scale invariance with those for conformal invariance for such heterotic sigma models, we find that the requirement for conformal invariance restricts further the geometry of $M^{D}$. In particular, there must exist a $\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ on $M^{D}$ such that $\widehat{V}=\theta+2d\Phi$, where $\theta$ is an appropriate Lee form that is constructed from the metric $G$ and the $\widehat{\nabla}$-covariantly constant forms on $M^{D}$ and $\Phi$ is the dilaton. If $\widehat{V}$ vanishes, then $\theta=-2d\Phi$ and the sigma model target space $M^{D}$ is conformally balanced. The conformal balance condition is an additional geometric restriction on $M^{D}$ – it usually arises as a condition for the solution of the dilatino Killing spinor equation of heterotic string effective theory. If $\widehat{V}$ does not vanish, the holonomy of the sigma model reduces to a subgroup of the groups stated above. In the latter case, we describe the pattern of holonomy reduction and investigate some aspects of the geometry. We also give examples of such manifolds based on a principal bundle construction – these involve an adaptation and in some cases a generalisation of constructions of Kähler and hyper-Kähler manifolds with torsion in poon ; Verbitsky .
In section two, we begin with a description of the invariance of the heterotic string effective action under reparameterisations of the spacetime $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations. After establishing this, we proceed with a Perelman style of argument to define the geometric $c$-function on the sigma model couplings and prove that it is decreasing along the renormalisation group flow of the heterotic sigma model. After that the proof of the statement that scale invariant heterotic sigma models with compact target spaces $M^{D}$ are conformally invariant follows as a consequence of the invariance of the geometric $c$-function under reparameterisations of $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations. In turn, this is a consequence of the invariance of the heterotic string effective action under these symmetries.
In section three, we demonstrate that sigma models with target spaces, whose the holonomy of the $\widehat{\nabla}$ connection is included in the groups listed in (1), are scale invariant. The proof requires a certain choice for the spacetime connection $\widetilde{\nabla}$ that contributes in the definition of $H$ in (5). Reviewing the work of HowePap ; deLaOssa , we point out that the heterotic sigma model action (2) is invariant under the symmetries generated by the $\widehat{\nabla}$-covariantly constant forms on manifolds with special holonomy groups that include those in (1). A special case of such a symmetry is that of $(0,p)$, $p=2,4$, worldsheet supersymmetry of heterotic sigma models previously described in Hullwitten – for completeness, we also present the off-shell superfield formulation of such heterotic heterotic sigma models given in HoweGP2 .
In section four, we point out that the additional requirement of conformal invariance for these models necessitates that the vector field $\widehat{V}=\theta+2{\mathrm{d}}\Phi$ to be $\widehat{\nabla}$-covariantly constant. There are two possibilities to consider. If $\widehat{V}=0$, then the target spaces are conformally balanced which is an additional restriction on $M^{D}$. The conformal balance condition together with a restriction on the holonomy of $\widehat{\nabla}$ imposes a strong restriction on the geometry of $M^{D}$. For example it is known that compact, conformally balanced Hermitian manifolds for which the holonomy of $\widehat{\nabla}$ is contained in $SU(n)$ and $H$ is closed are Calabi-Yau with $H=0$ SIGP3 . Alternatively, if $\widehat{V}\not=0$, the holonomy of $\widehat{\nabla}$ reduces further to a subgroup of the groups listed in (1). We identify the holonomy group reduction pattern and explore the local geometry of such sigma model target spaces. For this, we use the results of GLP , see also GGP for a review, where a similar holonomy reduction pattern has been observed in the context of classification of supersymmetric heterotic string backgrounds and its consequences on the geometry of a manifold have been explored. In particular, we find that the requirement of conformal invariance reduces the a priori $SU(n)$ holonomy to $\widehat{\nabla}$ to a subgroup of $SU(n-p)$ for $p=1,\dots,n-1$. Similarly, the a priori $Sp(k)$ holonomy of $\widehat{\nabla}$ reduces to $Sp(k-q)$ for $q=1,\dots,k$, while the a priori $G_{2}$ and $\mathrm{Spin}(7)$ holonomy of $\widehat{\nabla}$ reduce to a subgroup of $SU(3)$ and $G_{2}$, respectively. We also briefly explain how the $\widehat{\nabla}$-covariantly constant forms on $M^{D}$ associated with the holonomy groups (1) generate symmetries in the heterotic sigma model action (2).
2 Geometric $c$-theorem, scale and conformal invariance
2.1 Heterotic string effective action
The fields of a (0,1)-supersymmetric (heterotic) sigma-model consist of the maps $X$ from the superspace ${\mathbb{R}}^{2|1}$ with coordinates $(u,v|\vartheta^{+})$ into a Riemannian manifold $M^{D}$ with metric $G$ and the sections $\Psi$ of the vector bundle $S_{-}\otimes X^{*}E$, where $S_{-}$ is the anti-chiral spin bundle over ${\mathbb{R}}^{2|1}$ and $X^{*}E$ is the pull-back of a vector bundle over $M^{D}$ with $X$. Here $(u,v)$ are Grassmann even light-cone coordinates while $\vartheta^{+}$ is the Grassmann odd coordinate of ${\mathbb{R}}^{2|1,1}$. The action of an (0,1)-supersymmetric sigma-model with $G$, $B$ and $A$ couplings written in terms of (0,1) superfields $X$ Hullwitten is
$S=-{i\over 4\pi\alpha^{\prime}}\int_{{\mathbb{R}}^{2|1,1}}{\mathrm{d}}u{\mathrm{d}}v{\mathrm{d}}\vartheta^{+}\,\,\Big{(}(G+B)_{IJ}D_{+}X^{I}\partial_{v}X^{J}+ih_{ab}\Psi^{a}_{-}{\mathcal{D}}_{+}\Psi^{b}_{-}\Big{)}\,,$ (2)
where $D_{+}$ is a superspace derivative that commute with the supersymmetry generators and satisfy $D_{+}^{2}={\mathrm{i}}\partial_{u}$, and ${\mathcal{D}}_{+}\Psi^{a}_{-}=D_{+}\Psi_{-}^{a}+D_{+}X^{I}A_{I}{}^{a}{}_{b}\Psi_{-}^{b}$. Without loss of generality, we take the fibre metric $h$ of $E$ to satisfy ${\mathcal{D}}_{I}h_{ab}=0$. This action is manifestly invariant under (0,1) worldsheet supersymmetry transformations.
The classical action (2) of the heterotic sigma model is invariant under reparameterisations of $M^{D}$ and gauge transformations of $E$. These are generated by the infinitesimal transformations
$\delta X^{I}$ $=-V^{I}~{},~{}~{}~{}\delta G_{IJ}={\mathcal{L}}_{V}G_{IJ}~{},~{}~{}~{}\delta B_{IJ}={\mathcal{L}}_{V}B_{IJ}~{},~{}~{}~{}\delta A_{I}^{a}{}_{b}={\mathcal{L}}_{V}A_{I}^{a}{}_{b}\hfill$ (3)
$\delta\Psi_{-}^{a}$ $=-U^{a}{}_{b}\Psi_{-}^{b}~{},~{}~{}~{}\delta A_{I}^{a}{}_{b}={\mathcal{D}}_{I}U^{a}{}_{b}$ (4)
on the fields and couplings of the theory, where $V$ is a vector field on $M^{D}$ and $U$ is an infinitesimal gauge transformation of $E$. These transformations leave the action (2) invariant but they are not Noether style symmetries as they transform both the fields $X,\Psi$ and the couplings $G,B,A$ of the theory. So to avoid confusion with the usual Noether symmetries, they are referred to as sigma model symmetries. Introducing a frame for the metric $G$, $G_{IJ}=\delta_{AB}\mathbf{e}_{I}^{A}\mathbf{e}_{J}^{B}$, (2) will be invariant under frame rotations as well, see (7) below.
Although the classical action (2) is invariant under all these three sigma model symmetries, quantum mechanically all three are potentially anomalous moore ; Alvarez-Gaume ; Bagger . Such anomalies affect the geometric interpretation of the theory and as a consequence the invariance of the heterotic string effective theory under the above three transformations. As we have already been mentioned, the invariance of the heterotic string effective theory is essential for the Perelman style of argument for the $c$-theorem that we seek to establish. Because of this, we shall give a brief description how the effective heterotic string effective action is constructed in such a way that it is invariant under reparameterisation, frame and gauge transformations.
As it can always be arranged such that sigma model reparameterisation anomalies can be cancelled with the addition of a finite local counterterm in the sigma model effective action zumino , we shall focus from now on on the frame rotation and gauge transformation anomalies. The cancellation of these anomalies requires to assign an anomalous variation for $B$. This is done in such a way that
$H=dB+\frac{\alpha^{\prime}}{4}Q_{3}(\widetilde{\Omega},A)+{\cal O}(\alpha^{\prime 2})~{},$ (5)
is both frame rotations and gauge transformations invariant, where
$P_{4}(\widetilde{R},F)\equiv\widetilde{R}^{I}{}_{J}\wedge\widetilde{R}^{J}{}_{I}-F^{a}{}_{b}\wedge F^{b}{}_{a}\equiv\mathrm{tr}(\widetilde{R}\wedge\widetilde{R})-\mathrm{tr}(F\wedge F)=dQ_{3}(\widetilde{\Omega},A)~{},$ (6)
and $\widetilde{R}$ is the curvature of the frame connection $\widetilde{\Omega}$ on $M^{D}$. The choice of $\widetilde{\nabla}$ can be arbitrary as it can be replaced with any other connection on the spacetime up to the addition of a finite local counterterm in the sigma model the effective action. In sigma models with additional symmetry, like extended supersymmetry, a particular choice for $\widetilde{\nabla}$ has to be made for consistency. We shall elaborate in section 3 on this.
To give some more insight into (5), consider the frame rotations
$\delta\mathbf{e}^{A}_{I}=-\ell^{A}{}_{B}\,\mathbf{e}^{B}_{I}~{},~{}~{}~{}\delta_{\ell}\widetilde{\Omega}_{I}{}^{A}{}_{B}=\widetilde{\nabla}_{I}\ell^{A}{}_{B}~{},$ (7)
with infinitesimal parameter $\ell$. Varying the sigma model effective action $\Gamma$ with this transformation, one finds that it does not vanish. Instead, the variation gives an anomaly which can be expressed as
$\delta_{\ell}\Gamma=\frac{i\alpha^{\prime}}{16\pi}\int{\mathrm{d}}u{\mathrm{d}}v{\mathrm{d}}\vartheta^{+}\,\,Q_{2IJ}^{1}(\ell,\widetilde{\Omega})D_{+}X^{I}\partial_{v}X^{J}$ (8)
where $Q_{2}^{1}(\ell,\widetilde{\Omega})$ is constructed from the usual descent equations
$P_{4}(\widetilde{R})\equiv\mathrm{tr}(\widetilde{R}\wedge\widetilde{R})={\mathrm{d}}Q_{3}(\widetilde{\Omega})\Longrightarrow\delta_{\ell}P_{4}(\widetilde{R})=0={\mathrm{d}}\delta_{\ell}Q_{3}(\widetilde{\Omega})\Longrightarrow\delta_{\ell}Q_{3}={\mathrm{d}}Q_{2}^{1}(\ell,\widetilde{\Omega})~{},$ (9)
and $\widetilde{\Omega}$ is a frame connection with curvature $\widetilde{R}$.
To cancel the anomaly, one requires $B$ to attain an anomalous variation Hullwitten
$\delta_{\ell}B_{IJ}=-\frac{\alpha^{\prime}}{4}Q_{2IJ}^{1}(\ell,\widetilde{\Omega})~{},$ (10)
at first order in $\alpha^{\prime}$, i.e. in quantum theory at one loop in sigma model perturbation theory. Such an anomalous variation is part of the Green-Schwarz mechanism mgjs of cancelling reparameterisation, gauge and mixed anomalies in certain supergravity theories in ten dimensions. Noticing the sign difference in (10) and in the last term of (9), one can conclude that
$\delta_{\ell}H=0+{\cal O}(\alpha^{\prime 2})~{},$ (11)
at the indicated order in perturbation theory. Thus $H$ is invariant under frame rotations on $M^{D}$. A similar argument leads to the same conclusion for the invariance of $H$ in (5) under gauge transformations of $A$. Therefore, to construct a heterotic string effective action which is invariant under both frame rotations of $M^{D}$ and gauge transformations of $A$, one has to use $H$ as in (5) instead of simply writing $H={\mathrm{d}}B$. A direct consequence of setting $H$ as in (5) is that ${\mathrm{d}}H$ does not vanish, instead one has that
${\mathrm{d}}H=\frac{\alpha^{\prime}}{4}P_{4}(\widetilde{R},F)+{\cal O}(\alpha^{\prime 2})~{}.$ (12)
In the absence of other sources, the cancellation of the global anomaly also requires that $P_{4}(\widetilde{R},F)$ represents the trivial class in the cohomology of $M^{D}$ for $H$ to be globally defined. Furthermore, notice that the right hand side of the above equation vanishes provided that the vector bundle $E$ is identified with $TM^{D}$ and $F$ is chosen to be related to $\widetilde{R}$ so that the expression in (6) vanishes. This can only happen in special cases, like for example in sigma models with (1,1) worldsheet supersymmetry.
Having motivated the use of (5) for $H$ to construct the heterotic string effective action, let us turn to Euclidean signature backgrounds. Up to an overall constant that can be absorbed in an additive shift of $\Phi$, the standard action for the $G$, $B$, $A$ and $\Phi$ fields, in the order in $\alpha^{\prime}$ indicated [^4] From now on, the notation for the ${\cal O}(\alpha^{\prime 2})$ terms will be neglected. However, it will be always assumed that the formulae apply up to ${\cal O}(\alpha^{\prime 2})$ terms., is
$S(G,B,A,\Phi)=\int{\mathrm{d}}^{D}X\sqrt{G}\,e^{-2\Phi}\Big{(}-R-4G^{IJ}D_{I}\Phi D_{J}\Phi+\frac{1}{12}H^{2}$ (13)
$\qquad\qquad-\frac{\alpha^{\prime}}{8}\left(\widetilde{R}_{IJKL}\widetilde{R}^{IJKL}-F_{IJab}F^{IJab}\right)\Big{)}+{\cal O}(\alpha^{\prime 2})\,,$ (14)
where $H$ is given in (5) and $R$ is the Ricci scalar of $G$. From the discussion above, it is clear that the functional $S(G,B,\Phi)$ is invariant under the reparameterisations of $M^{D}$ and frame rotations of $M^{D}$, and the gauge transformations of $A$. In particular, for reparameterisation invariance $H$ must transform as a 3-form on $M^{D}$.
2.2 Geometric $c$-theorem and monotonicity of renormalization group flow
To prove a Perelman Perelman style $c$-theorem for heterotic sigma models, one begins with the heterotic string effective action (13) and rewrite it as follows:
$S(G,B,A,\Phi)=-\int{\mathrm{d}}^{D}X\sqrt{G}\,e^{-\Phi}\Big{(}-4\Delta^{2}+R-\frac{1}{12}H^{2}$ (15)
$\qquad\qquad+\frac{\alpha^{\prime}}{8}\left(\widetilde{R}_{IJKL}\widetilde{R}^{IJKL}-F_{IJab}F^{IJab}\right)\Big{)}e^{-\Phi}\,.$ (16)
Here $\Delta^{2}=G^{IJ}D_{I}D_{J}$ is the scalar Laplacian, where $D$ is the Levi-Civita connection of the metric $G$. Then, as it is explained elsewhere, one maximises $S(G,B,A,\Phi)$ as a function of $\Phi$ under the constraint
$C(G,\Phi)=\int{\mathrm{d}}^{D}X\sqrt{G}\,e^{-2\Phi}-1=0~{}.$ (17)
For this, it is sufficient to consider the eigenvalue problem
$\left(-4\Delta^{2}+R-\frac{1}{12}H^{2}+\frac{\alpha^{\prime}}{8}\left(\widetilde{R}_{IJKL}\widetilde{R}^{IJKL}-F_{IJab}F^{IJab}\right)\right)e^{-\Phi}=\lambda\,e^{-\Phi}\,,$ (18)
where $\lambda=\lambda(G,B,A)$ is the eigenvalue of the ground state. A Schrödinger-like operator has a unique normalized positive-definite eigenfunction associated with the ground state which can be identified with $e^{-\Phi}=e^{-\Phi}(G,B,A)$. This ground state is the unique extremum of $S(G,B,A,\Phi)$ under the constraint $C(G,\Phi)=0$, and in particular it is the absolute maximum as $S(G,B,A,\Phi)|_{\Phi=\Phi(G,B,A),C=0}=-\lambda(G,B,A)$.
Thus the constrained maximum of $S(G,B,\Phi)$ as a function of $\Phi$ for fixed $G$, $B$ and $A$ always exists and is unique. Because of this uniqueness, and general properties of elliptic partial differential equations, the maximum varies smoothly as a function of $G$, $B$ and $A$. Therefore, $\overline{S}(G,B,A)\equiv S(G,B,A,\Phi(G,B,A))$, where $\Phi$ is regarded as a function $\Phi(G,B,A)$, evaluated at the constrained maximum of $S$, is a smooth function of $G$, $B$ and $A$.
Next using that
$\delta\overline{S}=\Big{(}\frac{\delta\overline{S}}{\delta G_{IJ}}\delta G_{IJ}+\frac{\delta\overline{S}}{\delta B_{IJ}}\delta B_{IJ}+\frac{\delta\overline{S}}{\delta A_{Iab}}\delta A_{Iab}\Big{)}{\bigg{|}}_{C=0}$ (19)
$\qquad=\Big{(}\frac{\delta S}{\delta G_{IJ}}\delta G_{IJ}+\frac{\delta S}{\delta B_{IJ}}\delta B_{IJ}+\frac{\delta S}{\delta A_{Iab}}\delta A_{Iab}+\frac{\delta S}{\delta\Phi}\delta\Phi\Big{)}{\bigg{|}}_{\Phi=\Phi(G,B,A),C=0}~{},$ (20)
one finds that
$\frac{\delta\overline{S}}{\delta G_{IJ}}$ $=e^{-2\widehat{\Phi}}\beta_{G}^{IJ}=e^{-2\widehat{\Phi}}\Big{(}R^{IJ}-\frac{1}{4}H^{I}{}_{KL}H^{JKL}+2D^{I}D^{J}\Phi$ (21)
$+\frac{\alpha^{\prime}}{4}\left(\widetilde{R}^{I}{}_{KLN}\widetilde{R}^{JKLN}-F^{I}{}_{Kab}F^{JKab}\right)\Big{)}~{},$ (22)
$\frac{\delta\overline{S}}{\delta B_{IJ}}$ $=e^{-2\widehat{\Phi}}\beta_{B}^{IJ}=e^{-2\widehat{\Phi}}\left(\frac{1}{2}D_{K}H^{KIJ}-D_{K}\Phi H^{KIJ}\right)~{},$ (23)
$\frac{\delta\overline{S}}{\delta A_{Iab}}$ $=e^{-2\widehat{\Phi}}\beta_{A}^{Iab}=-\frac{\alpha^{\prime}}{2}\widehat{\nabla}_{J}\left(e^{-2\widehat{\Phi}}F^{JIab}\right)~{},$ (24)
where $\widehat{\Phi}=\Phi-\log\sqrt[{}^{4}]{G}$, $\beta_{G}$, $\beta_{B}$ and $\beta_{A}$ are the heterotic sigma model beta functions and ^$\widehat{\nabla}_{J}Y^{Ia}={\mathcal{D}}_{J}Y^{Ia}+\widehat{\Gamma}^{I}_{JK}Y^{Ka}$. Therefore from (21), one concludes that the first variation of the functional $\overline{S}$ along $G$, $B$ and $A$ can be written as
$\delta\overline{S}=\int{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\left(\beta_{G}^{IJ}\delta G_{IJ}+\beta_{B}^{IJ}\delta B_{IJ}+\beta_{A}^{Iab}\delta A_{Iab}\right)\,.$ (25)
One application of (25) is that $\overline{S}$ decreases along the heterotic sigma model renormalisation flow. Indeed, the renormalization group equation for the evolution of couplings $G$, $B$ and $A$ is
$\frac{{\mathrm{d}}G_{IJ}}{{\mathrm{d}}t}$ $=-\beta_{G,IJ}~{},~{}~{}~{}\frac{{\mathrm{d}}B_{IJ}}{{\mathrm{d}}t}=-\beta_{B,IJ}$ (26)
$\frac{{\mathrm{d}}A_{Iab}}{{\mathrm{d}}t}$ $=-\beta_{A,Iab}\,.$ (27)
Thus, one has that
$\frac{{\mathrm{d}}\overline{S}}{{\mathrm{d}}t}$ $=\int{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\left(\beta^{G,IJ}\frac{{\mathrm{d}}G_{IJ}}{{\mathrm{d}}t}+\beta^{B,IJ}\frac{{\mathrm{d}}B_{IJ}}{{\mathrm{d}}t}+\beta^{A,Iab}\frac{{\mathrm{d}}A_{Iab}}{{\mathrm{d}}t}\right)$ (28)
$=-\int{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\left(\beta^{G,IJ}\beta_{G,IJ}+\beta^{B,IJ}\beta_{B,IJ}+\beta^{A,Iab}\beta_{A,Iab}\right)\leq 0\,,$ (29)
This establishes a geometric version of the $c$-theorem up to order ${\cal O}(\alpha^{\prime 2})$ in sigma-model perturbation theory.
2.3 Conformal invariance from scale invariance
The condition for scale-invariance is not that $\beta_{G}$, $\beta_{B}$ and $\beta_{A}$ must vanish, but instead that they must vanish up to an infinitesimal diffeomorphism generated by a vector field $V$ on $M^{D}$ along with compensating $B$-field gauge transformation generated by a one-form $\Lambda$ and a gauge transformation of $A$ generated by $U$. Under the action of $V$, the variation of $G$, $B$ and $A$ is
$\delta G_{IJ}$ $=D_{I}V_{J}+D_{J}V_{I}\,,~{}~{}~{}\delta B_{IJ}=V^{K}H_{IJK}+\partial_{I}\Lambda_{J}-\partial_{J}\Lambda_{I}\,,$ (30)
$\delta A_{I}{}^{a}{}_{b}$ $=V^{J}F_{JI}{}^{a}{}_{b}+{\mathcal{D}}_{I}U^{a}{}_{b}\,,$ (31)
where $H$ in the variation for $B$ is given in (5). The presence of $H$ in the transformation of the $B$-field requires some explanation. This is because one may conclude that $B$ should transform with the Lie derivative along $V$ and so the transformation of $B$ should involve $dB$ instead of $H$ up to a $\Lambda$ transformation. However, we have seen that $B$ attains an anomalous variation in the quantum theory as a consequence of the anomaly cancellation mechanism for frame and gauge anomalies. As reparameterisation invariance is up to such transformations, this has to be taken into account in the transformation of $B$. For this one can set as the variation of $B$ to be
$\delta_{V}B_{IJ}={\mathcal{L}}_{V}B_{IJ}-\frac{\alpha^{\prime}}{4}Q_{2IJ}^{1}(\widetilde{\Omega},A,V)=V^{K}dB_{KIJ}-\frac{\alpha^{\prime}}{4}Q_{2IJ}^{1}(V,\widetilde{\Omega},A)+\partial_{I}\widetilde{\Lambda}_{J}-\partial_{J}\widetilde{\Lambda}_{I}~{},$ (32)
where the first term in the right hand side is expected for the transformation of $B$ under reparameterisations generated by $V$ while the second term in the right hand side is present to cancel the anomalies of the compensating frame and gauge transformations. The latter is computed from the descent equations
$P_{4}(\widetilde{\Omega},A)={\mathrm{d}}Q_{3}^{0}(\widetilde{\Omega},A)\Longrightarrow{\mathcal{L}}_{V}P_{4}(\widetilde{\Omega},A)=d{\mathcal{L}}_{V}Q_{3}^{0}(\widetilde{\Omega},A)\Longrightarrow$ (33)
$\iota_{V}P_{4}={\mathcal{L}}_{V}Q_{3}^{0}(\widetilde{\Omega},A)+{\mathrm{d}}Q_{2}^{1}(\widetilde{\Omega},A,V)~{},$ (34)
where the precise expression for $Q_{2}^{1}$ is not essential. Furthermore, from the last equation above, one concludes that
$Q_{2}^{1}(\widetilde{\Omega},A,V)=-\iota_{V}Q_{3}^{0}(\widetilde{\Omega},A)+{\mathrm{d}}M\,,$ (35)
for some 1-form $M$. Substituting this into (32)
$\delta_{V}B_{IJ}$ $=V^{K}dB_{KIJ}-\frac{\alpha^{\prime}}{4}Q_{2IJ}^{1}(V)+\partial_{I}\widetilde{\Lambda}_{J}-\partial_{J}\widetilde{\Lambda}_{I}$ (36)
$=(\iota_{V}dB)_{IJ}+\frac{\alpha^{\prime}}{4}(\iota_{V}Q_{3}^{0})_{IJ}+{\mathrm{d}}(\widetilde{\Lambda}-M)_{IJ}$ (37)
$=(\iota_{V}H)_{IJ}+{\mathrm{d}}\Lambda_{IJ}~{},$ (38)
as expected, where $\Lambda=\widetilde{\Lambda}-M$.
An alternative way to derive (38) is to recall that reparameterisation invariance of the effective action requires that $H$ transforms as a 3-form. Denoting the variation of $H$ under the diffeomorphisms generated by the vector field $V$ by $\overline{\delta}_{V}H$, one has
$\overline{\delta}_{V}H\equiv{\mathcal{L}}_{V}H\Longrightarrow{\mathrm{d}}\overline{\delta}_{V}B+\frac{\alpha^{\prime}}{4}\overline{\delta}_{V}Q^{0}_{3}(\widetilde{\Omega},A)=\frac{\alpha^{\prime}}{4}\iota_{V}P_{4}(\widetilde{\Omega},A)+{\mathrm{d}}\iota_{V}H\Longrightarrow$ (39)
${\mathrm{d}}\overline{\delta}_{V}B=\frac{\alpha^{\prime}}{4}\iota_{V}P_{4}(\widetilde{\Omega},A)-\frac{\alpha^{\prime}}{4}{\mathcal{L}}_{V}Q^{0}_{3}(\widetilde{\Omega},A)+{\mathrm{d}}\iota_{V}H\Longrightarrow$ (40)
${\mathrm{d}}\overline{\delta}_{V}B={\mathrm{d}}\iota_{V}H+\frac{\alpha^{\prime}}{4}{\mathrm{d}}Q_{2}^{1}(\widetilde{\Omega},A,V)$ (41)
where in the last step we have used (35). As a result, one has
$\overline{\delta}_{V}B=\iota_{V}H+\frac{\alpha^{\prime}}{4}Q_{2}^{1}(\widetilde{\Omega},A,V)+{\mathrm{d}}\Lambda~{},$ (42)
for some 1-form $\Lambda$. Taking into consideration the anomalous variation of $B$, one can set as the total variation of $B$,
$\delta_{V}B\equiv\overline{\delta}_{V}B-\frac{\alpha^{\prime}}{4}Q_{2}^{1}(\widetilde{\Omega},A,V)=\iota_{V}H+{\mathrm{d}}\Lambda~{},$ (43)
that reproduces (38) as expected.
Therefore, the conditions for scale invariance of the heterotic sigma model are
$B_{G\,IJ}$ $\equiv\beta_{G\,IJ}-(D_{I}V_{J}+D_{J}V_{I})=0\,,~{}~{}~{}~{}$ (44)
$B_{B\,IJ}$ $\equiv\beta_{B\,IJ}-(V^{K}H_{IJK}+\partial_{I}\Lambda_{J}-\partial_{J}\Lambda_{I})=0\,,~{}~{}~{}$ (45)
$B_{A\,Iab}$ $\equiv\beta_{A\,Iab}+\frac{\alpha^{\prime}}{2}\left(V^{J}F_{JI\,ab}+{\mathcal{D}}_{I}U_{ab}\right)=0\,,$ (46)
where the additional normalisation in the last condition is due to the additional normalisation of the beta function for $A$. The condition for local conformal invariance is instead $\beta_{G}=\beta_{B}=\beta_{A}=0$, in other words, the same condition but with $V=\Lambda=U=0$. By the Curci-Paffuti relation CP , if $\beta_{G}=\beta_{B}=\beta_{A}=0$, then $\beta_{\Phi}$ is a constant (the central charge) and conformal invariance holds, see also BGP for a geometric derivation of this up to order ${\cal O}(\alpha^{\prime 2})$.
Assuming that the sigma model is scale invariant $B_{G}=B_{B}=B_{A}=0$, one has that
00 0 $=\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\Big{(}\beta_{G}^{IJ}B_{G\,IJ}+\beta_{B}^{IJ}B_{B\,IJ}+\beta_{A}^{Iab}B_{A\,Iab}\Big{)}$ (47)
$=\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\left(\beta_{G}^{IJ}\beta_{G\,IJ}+\beta_{B}^{IJ}\beta_{B\,IJ}+\beta_{A}^{Iab}\beta_{A\,Iab}\right)\,,$ (48)
as the $\overline{S}$ is invariant under diffeomorphisms, $B$-field gauge transformations and the gauge transformations of $A$, i.e.
$\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\Big{(}\beta_{G}^{IJ}(D_{I}V_{J}+D_{J}V_{I})+\beta_{B}^{IJ}\left(V^{K}H_{IJK}+\partial_{I}\Lambda_{J}-\partial_{J}\Lambda_{I}\right)$ (49)
$\qquad\qquad-\frac{\alpha^{\prime}}{2}\beta_{A}^{Iab}\left(V^{J}F_{JI\,ab}+{\mathcal{D}}_{I}U_{ab}\right)\Big{)}=0\,,$ (50)
irrespective on whether the beta functions vanish. Therefore from (47), we conclude that $\beta_{G}=\beta_{B}=\beta_{A}=0$ and all scale invariant heterotic sigma models are conformally invariant with target spaces compact manifolds $M^{D}$. The proof above applies to up to two loops in sigma model perturbation theory. Using a similar argument as that presented in PapWitten for the the common sector of string theory, it can be extended to all loops in perturbation theory.
3 Geometry, scale and conformal Invariance
To give some examples of sigma models that exhibit scale and conformal invariance, we shall exploit some special geometries which naturally arise in the context of heterotic geometry. Typically, these involve an appropriate restriction on the holonomy of the connection $\widehat{\nabla}$ on $M^{D}$. It turns out that there are several such geometries that can be found. In all cases, the associated heterotic sigma models admit additional symmetries generated by the $\widehat{\nabla}$-covariantly constant forms of these geometries. These include additional worldsheet supersymmetries as well as additional W-type of symmetries. For the class of such geometries which will be explored below, the holonomy of $\widehat{\nabla}$ on $M^{D}$ is restricted to $U(n)$ $(D=2n)$, $SU(n)$ $(D=2n)$, $Sp(k)$ $(D=4k)$, $G_{2}$ $(D=7)$ and $\mathrm{Spin}(7)$ $(D=8)$, see also (1) in the introduction.
3.1 KT and CYT geometries
The Kähler with torsion (KT) and the Calabi-Yau with torsion (CYT) geometries [^5] KT, CYT and HKT (see section ) geometries are also referred to in the literature, especially for bi-KT, bi-CYT and bi-HKT, as generalised Kähler, generalised Calabi-Yau and generalised hyper-Kähler, respectively Hitchin ; Gualtieri – bi-KT, bi-CYT and bi-HKT geometries contain two copies of the associated structure one with respect to the connection $\widehat{\nabla}$ and the other with respect to the connection $\breve{\nabla}$. Here, we use the terminology of HoweGP to describe these geometries. The strong version of these geometries is also referred to as pluriclosed. have recently been reviewed PapWitten in a related context focusing on the strong case that the 3-form is closed, ${\mathrm{d}}H=0$. Because of this, here we shall only summarise without proof some of their key properties. However, we shall emphasise the differences that arise in the properties of these geometries whenever $H$ is not closed, ${\mathrm{d}}H\not=0$, as this is the case that applies to the heterotic sigma models, see (12). To begin suppose that $M^{D}$, $D=2n$, is a KT (Kähler with torsion) manifold, i.e. a Hermitian manifold equipped with the unique connection $\widehat{\nabla}$ with skew-symmetric torsion [^6] The torsion $H=-\iota_{I}{\mathrm{d}}I=-{\mathrm{d}}_{I}I$, where ${\mathrm{d}}_{I}=\iota_{I}{\mathrm{d}}-{\mathrm{d}}\iota_{I}={\mathrm{i}}(\partial-\overline{\partial})$, $\partial$ is the holomorphic exterior derivative and the inner derivation $\iota_{I}L$ of a $k$-form $L$ with respect to $I$ is $\iota_{I}L=\frac{1}{(k-1)!}I^{K}{}_{P_{1}}L_{KP_{2}\dots P_{k}}{\mathrm{d}}x^{P_{1}}\wedge\dots\wedge{\mathrm{d}}x^{P_{k}}$. $H$ such that both the metric $G$ and complex structure $I$ are covariantly constant, ^$\widehat{\nabla}G=\widehat{\nabla}I=0$. As consequence of ^$\widehat{\nabla}G=\widehat{\nabla}I=0$ is that the holonomy of $\widehat{\nabla}$ reduces from $SO(2n)$ to $U(n)$. The 3-form $H$ is of type [^7] On complex manifolds, a $k$-form can be decomposed in components characterised by a bi-degree $(r,s)$, with $r+s=k$, where $r$ denotes the number of holomorphic and $s$ denotes the number of anti-holomorphic directions of the form. $(2,1)\oplus(1,2)$ with respect to the complex structure $I$, i.e. its (3,0) and (0,3) components vanish. This is a consequence of ^$\widehat{\nabla}I=0$ together with the integrability of the complex structure $I$.
As $\widehat{\nabla}$ is a metric connection and ^$\widehat{\nabla}I=0$, the curvature tensor $\widehat{R}$ of the connection $\widehat{\nabla}$ is skew-symmetric in the last two indices, $\widehat{R}_{PQST}=-\widehat{R}_{PQTS}$, and satisfies the (1,1) condition
$\widehat{R}_{PQST}=\widehat{R}_{PQS^{\prime}T^{\prime}}\,I^{S^{\prime}}{}_{S}I^{T^{\prime}}{}_{T}\,.$ (51)
This condition is a consequence the reduction of the holonomy of $\widehat{\nabla}$ from $SO(2n)$ to $U(n)$ and arises as the integrability condition of ^$\widehat{\nabla}I=0$. The Lee form of a KT manifold is defined as
$\widehat{\theta}_{K}\equiv D^{P}I_{PQ}I^{Q}{}_{K}=-\frac{1}{2}I^{L}{}_{K}H_{LPQ}I^{PQ}\,,$ (52)
where the second equality follows as a consequence of ^$\widehat{\nabla}I=0$.
For the geometry on $M^{D}$ to satisfy the scale or conformal invariance condition, the KT structure has to be restricted further. This can be achieved by requiring that the holonomy of $\widehat{\nabla}$ is further reduced to $SU(n)\subset U(n)$ – such geometries are also referred to as CYT (Calabi-Yau with torsion) and have been previously investigated in the context of string compactifications in Strominger ; Hullheterotic . This in turn implies that the curvature $\widehat{R}$ satisfies the condition
$\widehat{R}_{PQST}\,I^{ST}=0\,,$ (53)
i.e. the Ricci form $\widehat{\rho}=\frac{1}{4}\widehat{R}_{PQST}\,I^{ST}dx^{P}\wedge dx^{Q}$ of the $\widehat{\nabla}$ connection vanishes. To demonstrate how this additional condition implies that the geometry on $M^{D}$ satisfies the scale invariance condition, consider that the Bianchi identity
$\widehat{R}_{K[LPQ]}=-\frac{1}{3}\widehat{\nabla}_{K}H_{LPQ}+\frac{1}{6}{\mathrm{d}}H_{KLPQ}\,.$ (54)
Notice that this contains the exterior derivative of $H$ as $H$ is not closed. Contracting this with the complex structure $I$, it yields
$\widehat{R}_{IJ}=\widehat{\nabla}_{I}\widehat{\theta}_{J}+\frac{1}{4}dH_{ILPQ}I^{L}{}_{J}I^{PQ}\,.$ (55)
after using both (51) and (53) as well as ^$\widehat{\nabla}I=0$ and the definition of the Lee form (52).
Further progress depends on being able to reexpress the $dH$ term in the equation above in terms of quadratic terms in $\widetilde{R}$ and $F$ that appear in the beta function for the metric in (21). One way to achieve this is to assume that $M^{D}$ admits a connection $\widetilde{\nabla}$ such that
$\widetilde{R}_{P^{\prime}Q^{\prime}KL}I^{P^{\prime}}{}_{P}I^{Q^{\prime}}{}_{Q}=\widetilde{R}_{PQKL}~{},~{}~{}~{}\widetilde{R}_{PQKL}I^{PQ}=0~{},$ (56)
and similarly for the connection [^8] These conditions (57) on a connection are a special case of Hermitian-Einstein or Hemitian-Yang-Mills equations and reduce on a 4-dimensional complex manifold to the usual anti-self-duality condition. Provided certain stability conditions are met, the Hermitian-Einstein equations have always a solution on any compact Kähler manifold as a consequence of Donalson-Uhlenbeck-Yau theorem Donaldson ; UhlenbeckYau and some classes of Hermitian manifolds LiYau , for a recent survey see Li . of vector bundle $E$
$F_{S^{\prime}T^{\prime}}{}^{a}{}_{b}I^{S^{\prime}}{}_{S}I^{T^{\prime}}{}_{T}=F_{ST}{}^{a}{}_{b}~{},~{}~{}~{}F_{PQ}{}^{a}{}_{b}I^{PQ}=0~{}.$ (57)
Thus both $\widetilde{R}$ and $F$ are (1,1) forms on $M^{D}$.
Using conditions (56) and (57), the equation (55) can be re-arranged as
$\widehat{R}_{IJ}+\frac{\alpha^{\prime}}{4}\left(\widetilde{R}_{IKLN}\widetilde{R}_{J}{}^{KLN}-F_{IKab}F_{J}{}^{Kab}\right)=\widehat{\nabla}_{I}\widehat{\theta}_{J}~{}.$ (58)
The symmetric and antisymmetric parts of this equation are actually equivalent to the conditions (44) for global scale-invariance, with $V$ and $\Lambda$ being related to $\widehat{\theta}$. A detailed investigation of the conditions will be described in section 4.1.
To the order in perturbation theory we consider, one can identify $\widetilde{R}$ with $\breve{R}$ – this is the curvature of the connection $\breve{\nabla}$ that has torsion $-H$. To justify this choice, the Bianchi identity that relates the curvatures of $\widehat{\nabla}$ and $\breve{\nabla}$ is
$\widehat{R}_{KLPQ}=\breve{R}_{PQKL}-\frac{1}{2}dH_{KLPQ}\,.$ (59)
Moreover, $dH$ in (55) is of order $\alpha^{\prime}$ as a consequence of (5) and (6). As $\widetilde{R}$ contributes in (55) at order $\alpha^{\prime}$, if we set $\widetilde{R}=\breve{R}$, the contribution of $H$ in $\breve{R}$ that is proportional to the Chern-Simons term $Q_{3}^{0}$ will be of order $\alpha^{\prime 2}$ in (55). This is an order higher than the one we are considering. Thus it suffices to set $\widetilde{R}=\breve{R}$, where $H$ in $\breve{R}$ is assumed to contribute at order zero in $\alpha^{\prime}$, i.e. $H=dB$. Thus the term $dH$ in (59) can be set to zero and so
$\widehat{R}_{KLPQ}=\breve{R}_{PQKL}\,.$ (60)
Thus at the order of $\alpha^{\prime}$ required, $\breve{R}$ satisfies the conditions (56) as a consequence of (51) and (53).
It remains to explore the condition for scale invariance associated with the gauge field coupling $A$ which is the last condition in (44). For this consider the Bianchi identity for $F$
${\mathcal{D}}_{K}F_{PQ}{}^{a}{}_{b}+{\mathcal{D}}_{Q}F_{KP}{}^{a}{}_{b}+{\mathcal{D}}_{P}F_{QK}{}^{a}{}_{b}=0$ (61)
and contract this with $I^{K}{}_{I}I^{PQ}$. After expressing ${\mathcal{D}}$ in terms of the connection with torsion, using ^$\widehat{\nabla}I=0$ and the conditions (57), one finds that
^$\widehat{\nabla}^{J}F_{JI}{}^{a}{}_{b}+\widehat{\theta}^{J}F_{JI}{}^{a}{}_{b}=0~{}.$ (62)
Clearly, the third condition for scale invariance in (44) is satisfied with $V$ related to $\widehat{\theta}$ and $U=0$. To conclude, if $M^{D}$ is a CYT manifold that admits connections $\widetilde{\nabla}$ and ${\mathcal{D}}$ that satisfy the conditions (56) and (57) and $H$ satisfies (12), then it solves the scale invariance conditions (44).
3.2 Scale invariance and other special geometries with torsion
3.2.1 HKT geometry
The Hyper-Kähler Geometry with Torsion (HKT) has also been reviewed in a similar context in PapWitten and so we shall only briefly explain the key points that will be useful later. An HKT manifold is a manifold of dimension $D=2n=4k$ endowed with three integrable complex structures $I_{r}$, $r=1,2,3$ that satisfy the algebra of imaginary unit quaternions: $I_{1}^{2}=I_{2}^{2}=-1$, $I_{1}I_{2}+I_{2}I_{1}=0$, and $I_{3}=I_{1}I_{2}$, or equivalently $I_{r}I_{s}=-\delta_{rs}1+\epsilon_{rs}{}^{t}I_{t}$. In addition, it is equipped with a metric $G$ and torsion $H$, which is a 3-form, such that the metric $G$ is Hermitian with respect to each of the complex structures $I_{r}$, $r=1,2,3$, and $I_{r}$ are covariantly constant, ^$\widehat{\nabla}I_{r}=0$, with respect to the connection $\widehat{\nabla}$. The geometry of the twistor space of an HKT manifolds and other properties have been investigated sometime ago in HoweGP . More recently, it has been shown that the moduli spaces of instantons on HKT manifolds admit an HKT structure RMMV following earlier work in LT ; NH . For a detailed description of the geometry of the moduli space of instantons on the HKT manifold $S^{3}\times S^{1}$ as well as its applications in AdS/CFT see Witten .
A key consequence of the definition is that the HKT manifold admits a KT structure with respect to each of the complex structures $I_{r}$, $r=1,2,3$. As a result, $H$ is a $(2,1)\oplus(1,2)$ form with respect to each of $I_{r}$, $r=1,2,3$ and $H=-{\mathrm{d}}_{I_{r}}I_{r}$, for each $r=1,2,3$. The latter condition, as it is valid for each value of $r$, imposes a restriction on the geometry. As in the KT case, $H$ is not necessarily a closed 3-form.
A consequence of the definition is that the three Lee forms, each associated to the three complex structures, of the HKT geometry are all equal
$\widehat{\theta}\equiv\widehat{\theta}_{1}=\widehat{\theta}_{2}=\widehat{\theta}_{3}~{}.$ (63)
This follows from the integrability of the complex structures and ^$\widehat{\nabla}I_{r}=0$. Also, it is valid irrespectively on whether $H$ is closed or not.
As $\widehat{\nabla}$ is a metric connection and ^$\widehat{\nabla}I_{r}=0$, the curvature $\widehat{R}$ of $\widehat{\nabla}$ is shew-symmetric in the last two indices and
$\widehat{R}_{PQST}=\widehat{R}_{PQS^{\prime}T^{\prime}}\,I_{r}^{S^{\prime}}{}_{S}I_{r}^{T^{\prime}}{}_{T}\,,~{}~{}~{}r=1,2,3~{},~{}~{}~{}(\mathrm{no~{}summation~{}over}~{}r)~{}.$ (64)
It turns out that this implies (53) for each of the complex structures $I_{r}$. The covariant constancy of $G$ and of the $I_{r}$ implies that the holonomy of $\widehat{\nabla}$ is contained in $Sp(k)\subset SU(n)\subset SO(D)$
Using the same arguments as in the CYT case, one can establish that the HKT geometry satisfies the scale invariant conditions (58) and (62) with Lee form given in (63) provided we impose the $Sp(k)$ instanton-like equations
$\widetilde{R}_{P^{\prime}Q^{\prime}KL}\,I_{r}{}^{P^{\prime}}{}_{P}\,I_{r}{}^{Q^{\prime}}{}_{Q}=\widetilde{R}_{PQKL}~{},~{}~{}~{}$ (65)
$F_{PQ}{}^{a}{}_{b}\,I_{r}{}^{P}{}_{K}\,I_{r}{}^{Q}{}_{L}=F_{KL}{}^{a}{}_{b}~{},~{}~{}~{}r=1,2,3~{},~{}~{}~{}(\mathrm{no~{}summation~{}over}~{}r)~{},$ (66)
on the curvature of $\widetilde{\nabla}$ and ${\mathcal{D}}$ and require that (12) is satisfied. The above conditions on $\widetilde{R}$ and $F$ are equivalent to the conditions that $\widetilde{R}$ and $F$ are $(1,1)$ forms with respect to each of the complex structures $I_{r}$. A similar argument to the KT case reveals that in perturbation theory to the order considered here, one can set $\widetilde{R}=\breve{R}$ in the expression for the scale invariance conditions (58).
3.2.2 $G_{2}$ and ${\mathrm{S}pin}(7)$ geometries with torsion
It is clear from the geometries investigated above that one way to construct solutions to the scale invariance conditions in (44) is to appropriately restrict the holonomy of $\widehat{\nabla}$. Apart from the two cases investigated so far, for which the holonomy of $\widehat{\nabla}$ is restricted to $SU(n)$ and $Sp(k)$, respectively, there are two more possibilities that can be explored. One is in dimension seven, $D=7$, for which the holonomy of $\widehat{\nabla}$ is restricted in $G_{2}\subset SO(7)$ and another in dimension eight, $D=8$, for which the holonomy of $\widehat{\nabla}$ is restricted in ${\mathrm{S}pin}(7)\subset SO(8)$.
To begin from the $G_{2}$ case, the $G_{2}$ geometry with torsion $H$, a 3-form, on $M^{7}$ with metric $G$ is characterised by the existence of a 3-form $\varphi$, the fundamental $G_{2}$ form, such that ^$\widehat{\nabla}\varphi=0$ tfsi1 ; tfsi2 , see also gmpw . In an adapted frame $({\mathbf{e}}^{A}={\mathbf{e}}^{i}_{I}dx^{I};A=1,\dots,7)$, where $x$ are some coordinates on the manifold, the metric and fundamental $G_{2}$ form $\varphi$ can be chosen as
$G=\delta_{AB}{\mathbf{e}}^{A}{\mathbf{e}}^{B}~{},~{}~{}~{}\varphi={\mathbf{e}}^{123}-{\mathbf{e}}^{167}-{\mathbf{e}}^{563}-{\mathbf{e}}^{527}-{\mathbf{e}}^{415}-{\mathbf{e}}^{426}-{\mathbf{e}}^{437}~{},$ (67)
where ${\mathbf{e}}^{123}$ is a shorthand for ${\mathbf{e}}^{1}\wedge{\mathbf{e}}^{2}\wedge{\mathbf{e}}^{3}$ and similarly for the rest of the components of $\varphi$. The condition [^9] The same condition ^$\widehat{\nabla}\varphi=0$ also determines $H$ in terms of $G$ and the fundamental form $\varphi$ and its first derivatives as for KT geometries. ^$\widehat{\nabla}\varphi=0$ implies that the curvature $\widehat{R}$ satisfies that
$\frac{1}{2}\widehat{R}_{PQK^{\prime}L^{\prime}}{}^{*}\varphi^{K^{\prime}L^{\prime}}{}_{KL}=\widehat{R}_{PQKL}~{},$ (68)
and $\widehat{R}_{PQKL}=-\widehat{R}_{PQLK}$ as $\widehat{\nabla}$ is metric, where ${}^{*}\varphi$ is the dual form of $\varphi$ on $M^{7}$. To see this, the integrability condition of ^$\widehat{\nabla}\varphi=0$ implies that
$\widehat{R}_{IJ}{}^{P}{}_{K}\varphi_{PLS}+\widehat{R}_{IJ}{}^{P}{}_{S}\varphi_{PKL}+\widehat{R}_{IJ}{}^{P}{}_{L}\varphi_{PSK}=0~{}.$ (69)
Contracting this with ${}^{*}\varphi^{KLS}{}_{T}$, one finds that
$\widehat{R}_{IJK^{\prime}L^{\prime}}\varphi^{K^{\prime}L^{\prime}}{}_{T}=0~{},$ (70)
where $\varphi^{IJ}{}_{K}{}^{*}\varphi_{IJPQ}=-4\varphi_{KPQ}$. Finally, contacting with $\varphi^{T}{}_{KL}$ and using
$\varphi^{QIJ}\varphi_{QKL}=(\delta^{I}_{K}\delta^{J}_{L}-\delta^{I}_{L}\delta^{J}_{K})-{}^{*}\varphi^{IJ}{}_{KL}~{},$ (71)
one derives (68).
Following the same reasoning as in the holonomy $SU(n)$ and $Sp(k)$ cases, one imposes the conditions
$\frac{1}{2}\widetilde{R}_{P^{\prime}Q^{\prime}KL}{}^{*}\varphi^{P^{\prime}Q^{\prime}}{}_{PQ}=\widetilde{R}_{PQKL}~{},~{}~{}~{}\frac{1}{2}F_{K^{\prime}L^{\prime}}{}^{a}{}_{b}{}^{*}\varphi^{K^{\prime}L^{\prime}}{}_{KL}=F_{KL}{}^{a}{}_{b}~{},$ (72)
on the curvatures $\widetilde{R}$ and $F$ of the connection $\widetilde{\nabla}$ and ${\mathcal{D}}$ – we shall refer to these equations as $G_{2}$ instanton-like conditions. Contracting the Bianchi identity (54) with ${}^{*}\varphi^{LPQ}{}_{T}$ and using the above equations, one confirms that the $G_{2}$ geometry with torsion satisfies the same scale invariance condition (58) as the KT geometry with Lee form $\widehat{\theta}$
$\widehat{\theta}_{L}\equiv-\frac{1}{6}D^{P}\varphi_{PSQ}\varphi^{SQ}{}_{L}=\frac{1}{6}H_{KPQ}{}^{*}\varphi^{KPQ}{}_{L}~{},$ (73)
where in the last step ^$\widehat{\nabla}\varphi=0$ has been used as well as (71). Again at the order in perturbation theory we are considering, one can set $\widetilde{R}=\breve{R}$ as in the KT case.
Furthermore, the $G_{2}$ geometry with torsion satisfies the same scale invariance condition for the gauge sector (62) as that of the CYT geometry with Lee form given in (73). The proof of this statement is similar to that given for CYT geometries and we shall not go into detail. Indeed, using the Bianchi identity for $F$ (61), appropriately contracting with ${}^{*}\varphi$, utilizing the $G_{2}$ instanton-like condition (72) and applying the identity
${}^{*}\varphi^{LIJK}{}^{*}\varphi_{LPQR}=6\,\delta^{I}_{[P}\delta^{J}_{Q}\delta^{K}_{R]}-9\,\delta^{[I}_{[P}\,{}^{*}\varphi^{JK]}{}_{QR]}~{},$ (74)
one can verify that $F$ satisfies (62).
Similarly to the $G_{2}$ geometry with torsion, the ${\mathrm{S}pin}(7)$ geometry with torsion $H$, a 3-form, on $M^{8}$ with metric $G$ is characterised by the existence of a self-dual 4-form $\phi$, the fundamental ${\mathrm{S}pin}(7)$ form, such that ^$\widehat{\nabla}\phi=0$ si2 . In an adapted frame $({\mathbf{e}}^{A}={\mathbf{e}}^{A}_{I}dx^{I};A=1,\dots,8)$ on $M^{8}$, the metric and fundamental form on $M^{8}$ can be written as
$G$ $=\delta_{AB}{\mathbf{e}}^{A}{\mathbf{e}}^{B}~{},~{}~{}~{}\phi={\mathbf{e}}^{1234}-{\mathbf{e}}^{1278}-{\mathbf{e}}^{1638}-{\mathbf{e}}^{1674}-{\mathbf{e}}^{5238}-{\mathbf{e}}^{5274}-{\mathbf{e}}^{5634}+{\mathbf{e}}^{5678}$ (75)
$\qquad\qquad-{\mathbf{e}}^{1526}-{\mathbf{e}}^{1537}-{\mathbf{e}}^{1548}-{\mathbf{e}}^{2637}-{\mathbf{e}}^{2648}-{\mathbf{e}}^{3748}~{}.$ (76)
The condition [^10] The same condition ^$\widehat{\nabla}\phi=0$ also determines $H$ in terms of $G$ and the fundamental form $\phi$ and its first derivatives as for KT geometries. on the curvature $\widehat{R}$ of $\widehat{\nabla}$ that arises as an integrability of ^$\widehat{\nabla}\phi=0$ can be expressed as
$\frac{1}{2}\widehat{R}_{PQK^{\prime}L^{\prime}}\phi^{K^{\prime}L^{\prime}}{}_{KL}=\widehat{R}_{PQKL}~{},$ (77)
where the curvature is skew-symmetric in the last two indices as $\widehat{\nabla}$ is metric. This can be shown following similar steps to those used to prove (68) and the identity
$\phi^{QPIJ}\phi_{QPKL}=6(\delta^{I}_{K}\delta^{J}_{L}-\delta^{I}_{L}\delta^{J}_{K})-4\phi^{IJ}{}_{KL}~{}.$ (78)
As in previous cases, we impose the conditions
$\frac{1}{2}\widetilde{R}_{P^{\prime}Q^{\prime}KL}\,\phi^{P^{\prime}Q^{\prime}}{}_{PQ}=\widetilde{R}_{PQKL}~{},~{}~{}~{}\frac{1}{2}F_{K^{\prime}L^{\prime}}{}^{a}{}_{b}\phi^{K^{\prime}L^{\prime}}{}_{KL}=F_{KL}{}^{a}{}_{b}~{},$ (79)
on the curvature of the $\widetilde{\nabla}$ connection and that of the gauge sector. Under these conditions one can again establish that this geometry satisfies the same scale invariance condition as that of the KT geometry (58) but now with Lee form
$\widehat{\theta}_{K}\equiv\frac{1}{36}D^{P}\phi_{PLSQ}\phi^{LSQ}{}_{K}=\frac{1}{6}H_{LSQ}\phi^{LSQ}{}_{K}~{}.$ (80)
To establish the second part of the identity one uses ^$\widehat{\nabla}\phi=0$ and (78). In the order of perturbation theory we are considering, one can again set $\widetilde{R}=\breve{R}$.
The gauge sector connections $A$ that satisfy the ${\mathrm{S}pin}(7)$ instanton-like condition (79) also satisfy the scale invariance condition of the Hermitian-Einstein connections (62) with the difference that Lee form is given in (80). The proof is similar to that outlined for the $G_{2}$ case above. Notice in particular that the fundamental form $\phi$ of the ${\mathrm{S}pin}(7)$ manifolds with torsion satisfy the identity (74) after replacing ${}^{*}\varphi$ with $\phi$ and appropriately adjusting the range of the indices.
3.3 Worldsheet symmetry and geometry
$\widehat{\nabla}$-covariantly constant forms on the spacetime $M^{D}$ generate symmetries in sigma model actions, like for example that of (2) HowePap . More specifically, if the sigma model target space $M^{D}$ admits a $\widehat{\nabla}$ covariantly constant $(\ell+1)$-form $L$, i.e.
^$\widehat{\nabla}_{I}L_{J_{1}\dots J_{\ell+1}}=0~{},$ (81)
then the action (2) is invariant under the transformations
$\delta_{L}X^{I}=\varepsilon_{L}L^{I}{}_{J_{1}\dots J_{\ell}}D_{+}X^{J_{1}}\cdots D_{+}X^{J_{\ell}}~{},~{}~{}~{}\delta_{L}\Psi_{-}^{a}=-A_{I}{}^{a}{}_{b}\delta_{L}X^{I}\Psi_{-}^{b}~{},$ (82)
provided that in addition the curvature $F$ of the gauge sector satisfies deLaOssa
$F_{K[J_{1}}L^{K}{}_{J_{2}\dots J_{\ell+1}]}=0~{},$ (83)
where $\varepsilon_{L}$ is the infinitesimal parameter whose Grassmannian parity is chosen such that the variation $\delta_{L}X$ is even. It turns out that the conditions imposed on $F$ in (57), (66), (72) and (79) all follow from the condition (83) that arises from the requirement of invariance of heterotic sigma model action under the transformations (82) generated by the fundamental forms of the holonomy groups $SU(n)$, $Sp(k)$, $G_{2}$ and $\mathrm{Spin}(7)$ of $\widehat{\nabla}$ connection. This justifies the restrictions we have put on the curvature of the gauge sector in the discussion of geometries that satisfy the scale invariance conditions.
The commutator, $[\delta_{L},\delta_{P}]X$ of two such transformations on $X$ generated by the $(\ell+1)$-form $L$ and the $(p+1)$-form $P$ on $M$ is rather involved to be described here, see HPS for a detailed exposition. Even in the $G_{2}$ and ${\mathrm{S}pin}(7)$ cases, where the symmetries are generated by the fundamental forms that exhibit special algebraic properties, the expressions are rather involved especially in the case that $H\not=0$. In the $G_{2}$ and ${\mathrm{S}pin}(7)$, the algebra closes as a W-algebra, i.e. the structure constants of the algebra depend on conserved currents. The commutator on $\Psi$ is
$[\delta_{L},\delta_{P}]\Psi_{-}^{a}=-A_{K}{}^{a}{}_{b}[\delta_{L},\delta_{P}]X^{K}-F_{IJ}{}^{a}{}_{b}\delta_{L}X^{I}\delta_{P}X^{J}\Psi_{-}^{b}~{}.$ (84)
In all cases of interest here, the second term in the commutator can be determined using the invariance condition (83). Therefore, the closure properties of the transformations on $\Psi$ are essentially dictated by those on $X$.
The simplest cases to consider are those for which the sigma-model target space admits either a KT structure $I$ (^$\widehat{\nabla}I=0$), or an HKT structure $I_{r}$ (^$\widehat{\nabla}I_{r}=0$), then the action (2) is invariant under (0,2) or (0,4) worldsheet supersymmetry transformations [^11] There is an extensive literature, see for example Curtrightfreedman ; LAGDF ; Curtrightzachos ; Howesierra ; JGCHMR ; Buscher ; Braaten ; Hullreview ; HoweGP1 , on the geometry of target spaces of 2-dimensional sigma models with $(p,q)$, $p,q\not=0$, supersymmetry following the early work of Zumino2 in four dimensions., respectively. The additional supersymmetry transformations Hullwitten are given by
$\delta_{I_{r}}X^{K}=\varepsilon^{r}I_{r}^{K}{}_{L}D_{+}X^{L}\,,~{}~{}~{}\delta_{I_{r}}\Psi_{-}^{a}=-A_{K}{}^{a}{}_{b}\delta_{I_{r}}X^{K}\Psi_{-}^{b}$ (85)
with $r=1$ for the KT case and $r=1,2,3$ for the HKT case, where $\varepsilon^{r}=\varepsilon^{r}(u,\vartheta^{+})$ are Grassmannian odd infinitesimal parameters. These can depend on the worldsheet superspace coordinates $(u,\vartheta^{+})$, as classically the action is superconformally invariant. The commutator of two such transformations on $X$ and $\Psi$ are given by
$[\delta_{I},\delta^{\prime}_{I}]X^{K}=-2{\mathrm{i}}\varepsilon^{r}\varepsilon^{\prime s}\delta_{rs}\partial_{u}X^{K}+(\varepsilon^{\prime s}D_{+}\varepsilon^{r}I_{s}^{K}{}_{P}I_{r}^{P}{}_{L}-\varepsilon^{r}D_{+}\varepsilon^{\prime s}I^{K}_{r}{}_{P}I^{P}_{s}{}_{L})D_{+}X^{L}\,,$ (86)
$[\delta_{I},\delta^{\prime}_{I}]\Psi^{a}=-A_{K}{}^{a}{}_{b}[\delta_{I},\delta^{\prime}_{I}]X^{K}+\varepsilon^{r}\,\varepsilon^{\prime s}\delta_{rs}F_{IJ}{}^{a}{}_{b}\,D_{+}X^{I}\,D_{+}X^{J}\,\Psi^{b}_{-}~{},$ (87)
where the integrability condition of the complex structures $I_{r}$ has been used [^12] Vanishing of the Nijenhuis tensor of $I_{r}$ simplifies the right hand side of the commutator (87). together with the conditions (83) for $L=I_{r}$. The latter conditions imply that $F$ is a (1,1)-form with respect to each complex structure $I_{r}$. In the KT case, the second term in the right hand side of the commutator can be expressed as $(\varepsilon D_{+}\varepsilon^{\prime}-\varepsilon^{\prime}D_{+}\varepsilon)D_{+}X^{K}$, while in the HKT case it can be expressed as $\delta_{rs}(\varepsilon^{r}D_{+}\varepsilon^{\prime s}-\varepsilon^{\prime s}D_{+}\varepsilon^{r})D_{+}X^{K}-(\varepsilon^{r}D_{+}\varepsilon^{\prime s}+\varepsilon^{\prime s}D_{+}\varepsilon^{r})\epsilon_{rs}{}^{t}I_{t}{}^{K}{}_{L}D_{+}X^{L}$. In either case, the commutator (87) closes to spacetime translations and supersymmetry transformations as expected. The commutator of two transformations (85) on $\Psi$ closes on-shell to the standard $(0,q)$ supersymmetry algebra, for either $q=2$ or $q=4$.
[…]
$D_{+r}X^{K}=I^{K}_{r}{}_{L}D_{+0}X^{L}~{},~{}~{}~{}{\mathcal{D}}_{+r}\Psi_{-}^{a}=J^{a}_{r}{}_{b}{}{\mathcal{D}}_{+0}\Psi_{-}^{b}~{},$ (88)
where $J_{r}$ are fibre complex structures on $E$ with $r=1$ for $(0,2)$ supersymmetry and $r=1,2,3$ for (0,4) supersymmetry. The integrability conditions of these constraints on $X$ yield all the conditions that arise from the closure of the algebra of $(0,q)$ supersymmetry transformations on $X$. There are additional conditions that arise as the integrability conditions of the constraints on $\Psi$. These make a rather long list which can be found in HoweGP2 . It turns out that for (0,2) supersymmetry is sufficient to require that $J=J_{1}$ is a fibre complex structure, ${\cal D}_{I}J^{a}{}_{b}=0$ and that $F$ is a (1,1)-form with respect to $I$ on $M^{D}$ $(D=2n)$. For (0.4) supersymmetry, it is required that $J_{r}$ satisfy the algebra of imaginary unit quaternions, ${\cal D}_{I}J_{r}{}^{a}{}_{b}=0$ and that $F$ is a (1,1)-form on $M^{D}$, $(D=4k)$, with respect to all complex structures $I_{r}$. An action for these multiplets is
$S=-{i\over 4\pi\alpha^{\prime}}\int_{{\mathbb{R}}^{2|1,1}}{\mathrm{d}}u{\mathrm{d}}v{\mathrm{d}}\vartheta^{0}\,\,\Big{(}(G+B)_{IJ}D_{+0}X^{I}\partial_{v}X^{J}+ih_{ab}\Psi_{-}^{a}{\mathcal{D}}_{+0}\Psi_{-}^{b}\Big{)}\,,$ (89)
where the superfields $X$ and $\Psi$ satisfy the constraint (88). It can be shown that this action is invariant under all $(0,q)$ supersymmetry transformations provided that in addition the fibre metric $h$ is (1,1) tensor with respect to the fibre complex structures $J_{r}$, i.e. $h_{a^{\prime}b^{\prime}}J_{r}{}^{a^{\prime}}{}_{a}J_{r}{}^{b^{\prime}}{}_{b}=h_{ab}$ for either $r=1$ or $r=1,2,3$.
4 Scale and conformal invariance revisited
4.1 Comparison of conditions for scale and conformal invariance
We have shown that all special geometries we have investigated in section 3 satisfy up to order ${\cal O}(\alpha^{\prime 2})$ the scale invariance conditions
$\widehat{R}_{IJ}+\frac{\alpha^{\prime}}{4}\left(\widetilde{R}_{IKLN}\widetilde{R}_{J}{}^{KLN}-F_{IKab}F_{J}{}^{Kab}\right)=\widehat{\nabla}_{I}\widehat{\theta}_{J}~{},$ (90)
^$\widehat{\nabla}^{J}F_{JI}{}^{a}{}_{b}+\widehat{\theta}^{J}F_{JI}{}^{a}{}_{b}=0~{},$ (91)
provided that (12) holds and the connection $\widetilde{\nabla}$ and ${\mathcal{D}}$ satisfy instanton-like conditions. A comparison of these conditions for scale invariance with the conformal invariance conditions, which are the vanishing conditions for the beta functions of $G$, $B$ and $A$, in (21), reveals that these special geometries solve the latter provided that the 1-form
$\widehat{V}_{L}=\widehat{\theta}_{L}+2D_{L}\Phi\,,$ (92)
must satisfy
^$\widehat{\nabla}_{K}\widehat{V}_{L}=0~{},~{}~{}~{}\widehat{V}^{K}F_{KL}{}^{a}{}_{b}=0\,,$ (93)
where $\Phi$ is the dilaton. Therefore, $\widehat{V}$ is $\widehat{\nabla}$-covariantly constant and the directions of $F$ along $\widehat{V}$ vanish. The former condition implies that
${\mathcal{L}}_{\widehat{V}}G=0~{},~{}~{}~{}{\mathrm{d}}\widehat{V}=\iota_{\widehat{V}}H\,,$ (94)
i.e. $\widehat{V}$ is a Killing vector field.
It is clear that from the conditions (93) that conformal invariance restricts further the special geometry on $M^{D}$. There are two distinct possibilities to be considered depending on whether $\widehat{V}$ vanishes or not leading to two different types of additional geometric structure on $M^{D}$. If $\widehat{V}$ vanishes, $\widehat{V}=0$, this trivially solves the conditions for conformal invariance (93) and expresses the Lee form of the special geometries in terms of the dilaton, $\widehat{\theta}=-2{\mathrm{d}}\Phi$. Geometries with this property are called conformally balanced [^13] This terminology has been adapted from differential geometry where a manifold is balanced, iff the Lee form of a fundamental form the holonomy group vanishes. and they can be restrictive. For example, compact conformally balanced CYT manifolds with $dH=0$, or with $dH\not=0$ but otherwise appropriately restricted, can be shown to be Calabi-Yau with $H=0$ SIGP3 . Conformally balanced geometries have arisen before in the context of (spacetime) supersymmetric heterotic string backgrounds and this relation will be explored in more detail in the next section.
Alternatively, suppose that $\widehat{V}\not=0$. As $\widehat{V}$ is non-vanishing and $\widehat{\nabla}$-covariantly constant, it is non-vanishing everywhere on $M^{D}$. Moreover its length is constant as $\widehat{\nabla}$ is metric. The integrability condition of ^$\widehat{\nabla}\widehat{V}=0$ reveals that
$\widehat{R}_{IJKL}\widehat{V}^{L}=0~{},$ (95)
and the holonomy of the connection $\widehat{\nabla}$ reduces further to a subgroup of the original holonomy group stated in (1) whose geometry was summarised in sections 3.1 and 3.2. The pattern of these holonomy reductions depends on the original holonomy group and they will be explored below together with some aspects of the geometry of $M^{D}$. Here, we shall proceed to show that the couplings are invariant under the diffeomorphisms generated by $\widehat{V}$.
We have already demonstrated that $\widehat{V}$ is Killing and so the metric $G$ is invariant under the diffeomorphisms generated by $\widehat{V}$. If either $\widetilde{\nabla}$ is chosen such
$\widehat{V}^{K}\widetilde{R}_{KLIJ}=0~{},$ (96)
or at the order in the perturbation theory considered here one sets $\widetilde{R}=\breve{R}$ in (12), then
${\mathcal{L}}_{\widehat{V}}H={\mathrm{d}}\iota_{\widehat{V}}H+\iota_{\widehat{V}}{\mathrm{d}}H={\mathrm{d}}^{2}\widehat{V}+\frac{\alpha^{\prime}}{4}\iota_{\widehat{V}}P_{4}(\widetilde{R},F)=0~{},$ (97)
and $H$ is invariant as well.
The curvature $F$ of $A$ is invariant under the transformations generated by $\widehat{V}$ up to a gauge transformation. Indeed
${\mathcal{L}}_{\widehat{V}}F=\iota_{\widehat{V}}dF=\iota_{\widehat{V}}{\mathcal{D}}F-\iota_{\widehat{V}}([A,F])=-[\iota_{\widehat{V}}A,F]~{},$ (98)
where we have used the Bianchi identity for $F$ and the second equation in (93). As a consequence, under the stated assumptions, all the couplings constants of the sigma model are invariant under the diffeomorphisms generated by $\widehat{V}$.
4.2 Geometry of conformally invariant backgrounds with $\widehat{V}=0$ and spacetime supersymmetry
We have demonstrated that the special geometries investigated in section 3 that satisfy the conformal invariance conditions with $\widehat{V}=0$ are conformally balanced, i.e. $\widehat{\theta}=-2{\mathrm{d}}\Phi$. The conformal balance condition has arisen before in the context of (spacetime) supersymmetric heterotic backgrounds. Here, we shall address the question whether the invariance conditions (81) and (83) on the sigma model couplings we have found in section 3 together with the conformal balanced condition on $M^{D}$ imply that the geometry of $M^{D}$ solves the heterotic supergravity Killing spinor equations (KSEs), i.e. $M^{D}$ is a (spacetime) supersymmetric background.
In all cases, the gravitino KSE admits non-trivial solutions as a consequence of the restriction of the holonomy group of the $\widehat{\nabla}$ connection to one of those in (1). The same applies for the gaugino Killing spinor equation as a consequence of the condition (83) on the gauge coupling of the sigma model that arises from the requirement of the invariance of the sigma model action under the symmetries generated by the $\widehat{\nabla}$-covariantly constant forms associated with the holonomy group. Furthermore, the conformal balance condition always arises as one of the conditions of the dilatino KSE. So the target space of a conformally invariant sigma model whose couplings satisfy (81) and (83) is a supersymmetric background provided that the dilatino KSE does not impose additional conditions on the geometry of $M^{D}$ to that of conformal balance. To explore this statement requires a case by case investigation.
First consider the CYT geometries for which the holonomy of $\widehat{\nabla}$ is contained in $SU(n)$. The dilatino KSE for such geometries implies that $M^{2n}$ is a conformally balanced complex manifold. As the integrability of the complex structure required by the dilatino KSE is already included in the definition of CYT structure, the target space of all conformally invariant sigma models with $\widehat{V}=0$ whose couplings satisfy the conditions required for the invariance of the sigma model action under the transformations generated by the $\widehat{\nabla}$-covariantly constant forms are (spacetime) supersymmetric. A similar conclusion holds for conformally balanced HKT and $\mathrm{Spin}(7)$ geometries.
Next let us turn to the $G_{2}$ case. The dilatino KSE, apart from the conformal balance condition on $M^{7}$, it also imposes the condition that $H_{IJK}\varphi^{IJK}=0$. It is clear from the use of the Bianchi identity (54), the condition that the holonomy of $\widehat{\nabla}$ is included in $G_{2}$, the expression for ${\mathrm{d}}H$ in (12) and (72) that $H_{IJK}\varphi^{IJK}$ must be constant. The dilatino KSE implies that this constant vanish. This is an additional condition on the geometry which does not arise from either the conformal invariance of the sigma model or the invariance of the sigma model action under the symmetries generated by the $\widehat{\nabla}$-covariantly constant $G_{2}$ forms. Therefore, conformal invariance and the invariance of the sigma model actions under the symmetries generated by the $\widehat{\nabla}$-covariantly constant $G_{2}$ forms does not necessarily imply that the sigma model target space is supersymmetric. Though, we have not been able to provide an example to demonstrate that this obstruction to spacetime supersymmetry occurs.
4.3 Geometry of conformally invariant backgrounds with $\widehat{V}\not=0$
As $\widehat{V}$ is non-vanishing and $\widehat{\nabla}$-covariantly constant, the holonomy of $\widehat{\nabla}$ reduces further to a subgroup of the groups stated in (1). If the a priori holonomy group of $\widehat{\nabla}$ is one of the groups listed in (1), this subgroup will be specified as the invariance group of all $\widehat{\nabla}$-covariantly constant forms on $M^{D}$, including $\widehat{V}$. The pattern of holonomy reduction depends on the original a priori holonomy group of $\widehat{\nabla}$ and it will be outlined below in each case.
4.3.1 KT and CYT geometries and conformal invariance
Suppose that $M^{D}$ is a KT (or CYT) manifold that satisfies the conditions for conformal invariance with $\widehat{V}\not=0$. As it has already been mentioned the length of $\widehat{V}$ is constant, it is convenient in what follows to normalise $\widehat{V}$ to have length $1$,
$G(\widehat{V},\widehat{V})=1~{},$ (99)
after possibly re-scaling it with an appropriate constant. As ^$\widehat{\nabla}\widehat{V}=0$ and ^$\widehat{\nabla}I=0$, one has that
^$\widehat{\nabla}\widehat{Z}=0~{},~{}~{}~{}\widehat{Z}^{J}=-I^{J}{}_{K}\widehat{V}^{K}~{}.$ (100)
The vector fields $\widehat{V}$ and $\widehat{Z}$ are orthogonal $G(\widehat{V},\widehat{Z})=0$ and $\widehat{Z}$ has length $1$ as well. Furthermore, $\widehat{Z}$, as $\widehat{V}$, leaves invariant $G$, $H$ and $F$ – the latter up to a gauge transformation. Indeed $\widehat{Z}$ is Killing as ^$\widehat{\nabla}\widehat{Z}=0$ and
${\mathcal{L}}_{\widehat{Z}}H={\mathrm{d}}\iota_{\widehat{Z}}H+\iota_{\widehat{Z}}{\mathrm{d}}H={\mathrm{d}}^{2}\widehat{Z}+\frac{\alpha^{\prime}}{4}\iota_{\widehat{Z}}P_{4}(\widetilde{R},F)=0~{},$ (101)
where we have used $\iota_{\widehat{Z}}F=0$ that in turn follows from $\iota_{\widehat{V}}F=0$ and that $F$ is (1,1)-form with respect to $I$. We have also taken $\iota_{\widehat{Z}}\widetilde{R}=0$ as we have done for $\widehat{V}$ in (96). The invariance of $F$ follows from $\iota_{\widehat{Z}}F=0$ and the argument stated in (98). One consequence of ^$\widehat{\nabla}\widehat{V}=\widehat{\nabla}\widehat{Z}=0$ is that the holonomy of $\widehat{\nabla}$ reduces to a subgroup of $U(n-1)$ in the KT case and to a subgroup of $SU(n-1)$ in the CYT case.
To continue, let us assume that the $[\widehat{V},\widehat{Z}]$ does not vanish. Using ${\mathrm{d}}\widehat{V}=\iota_{\widehat{V}}H$ and similarly for $\widehat{Z}$, one has
$[\widehat{V},\widehat{Z}]=\iota_{\widehat{V}}\iota_{\widehat{Z}}H~{},$ (102)
where it has been used that $G(\widehat{V},\widehat{Z})=0$. Therefore, the commutator of the two vector fields is given by components of $H$. As a consequence of this, the Bianchi identity (54), the condition (96) on $\widetilde{R}$ for both $\widehat{V}$ and $\widehat{Z}$ and (12), the commutator $[\widehat{V},\widehat{Z}]$ of $\widehat{V}$ with $\widehat{Z}$ is also a $\widehat{\nabla}$-covariantly constant vector field,
^$\widehat{\nabla}[\widehat{V},\widehat{Z}]=0~{}.$ (103)
As a result $[\widehat{V},\widehat{Z}]$ is Killing. It can also be shown that $\iota_{[\widehat{V},\widehat{Z}]}F=0$. Indeed,
$\iota_{[\widehat{V},\widehat{Z}]}F^{a}_{I}$ $=[\widehat{V},\widehat{Z}]^{J}F^{a}_{JI}=(\widehat{V}^{L}D_{L}\widehat{Z}^{J}-\widehat{Z}^{L}D_{L}\widehat{V}^{J})F_{JI}^{a}=\widehat{V}^{L}\widehat{Z}^{J}(\mathcal{D}_{L}F^{a}_{JI}-\mathcal{D}_{J}F^{a}_{LI})$ (104)
$=-\widehat{V}^{L}\widehat{Z}^{J}\mathcal{D}_{I}F^{a}_{LJ}=D_{I}(\widehat{V}^{L}\widehat{Z}^{J})F^{a}_{LJ}=(D_{I}\widehat{V}^{L}\widehat{Z}^{J}+\widehat{V}^{L}D_{I}\widehat{Z}^{J})F^{a}_{LJ}=0~{},$ (105)
where we have used that $\iota_{\widehat{V}}F=\iota_{\widehat{Z}}F=0$ as well as the Bianchi identity for $F$. Consequently, $[\widehat{V},\widehat{Z}]$ leaves invariant both $H$ and $F$ – the latter up to a gauge transformation. The arguments for these are similar to those stated in (96) and (101), respectively. Furthermore, $[\widehat{V},\widehat{Z}]$ is orthogonal to both $\widehat{V}$ and $\widehat{Z}$, i.e.
$G([\widehat{V},\widehat{Z}],\widehat{V})=G([\widehat{V},\widehat{Z}],\widehat{Z})=0~{},$ (106)
and
$I([\widehat{V},\widehat{Z}],\widehat{V})=I([\widehat{V},\widehat{Z}],\widehat{Z})=0~{}.$ (107)
The above two equations can be demonstrated using that ^$\widehat{\nabla}G=\widehat{\nabla}I=0$ and (102) as well as the definition of $\widehat{Z}$. One consequence of the above relations is that after normalizing $[\widehat{V},\widehat{Z}]$ to have length $1$, as it was done for $\widehat{V}$, one can repeat the same process as that we have demonstrated with $\widehat{V}$ but now with $[\widehat{V},\widehat{Z}]$ in its place in the directions orthogonal to $\widehat{V}$ and $\widehat{Z}$. This means to consider the vector field $-I([\widehat{V},\widehat{Z}])$, which is again $\widehat{\nabla}$ covariantly constant, and observe that leaves invariant [^14] We also require that the condition (96) on $\widetilde{R}$ is satisfied with $\widehat{V}$ replaced by $[\widehat{V},\widehat{Z}]$. This is justified as $\breve{R}$ satisfies the same condition at the order of the perturbation theory considered. $G$, $H$ and $F$. This can be continued until the algebra of $\widehat{\nabla}$-covariantly constant vector fields closes to a Lie algebra $\mathfrak{g}$ of dimension $2p$. In such a case, the holonomy of $\widehat{\nabla}$ will reduce to $U(n-p)$ for KT geometries and to $SU(n-p)$ for CYT geometries.
As the vector fields generated by the action of $\mathfrak{g}$ on $M^{2n}$ have no fixed points, because they are nowhere vanishing, the action of $\mathfrak{g}$ on $M^{2n}$ is free. This can be integrated to an effective action of the (unique) group $\mathcal{G}$ on $M^{2n}$ which is the universal cover of all groups with Lie algebra $\mathfrak{g}$. The action of $\mathcal{G}$ on $M^{2n}$ is almost free, i.e. the isotropy group of any fixed points is a discrete subgroup of $\mathcal{G}$. This geometry is reminiscent of that encountered in the investigation of supersymmetric heterotic backgrounds in GLP and as a result we shall adopt a similar language to describe it. Assuming that $\mathcal{G}$, or at least an appropriate quotient of $\mathcal{G}$ with a discrete subgroup, acts freely (on the space of principal orbits) of $M^{2n}$, the geometry of $M^{2n}$ can be organised as follows: Let $(\lambda^{\alpha};a=1,\dots,2p)$ be the 1-forms dual to the $\widehat{\nabla}$-covariantly constant vector fields $(\widehat{V}_{\alpha};a=1,\dots,2p)$ generated by the action of $\mathfrak{g}$ on $M^{2n}$. One can view $(\lambda^{\alpha};\alpha=1,\dots,2p)$ as a principal bundle connection, $\lambda^{\alpha}(\widehat{V}_{\beta})=\delta^{\alpha}_{\beta}$. As $\lambda^{\alpha}$ is no-where vanishing and $G^{-1}(\lambda^{\alpha},\lambda^{\beta})=\delta^{\alpha\beta}$, because we have normalised $\widehat{V}_{\alpha}$ to have length $1$, it can be used to construct an orthonormal (co-)frame on $M^{2n}$ as $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$. Then, the metric $G$ and $H$ can be written as
$G$ $\equiv\delta_{\alpha\beta}\lambda^{\alpha}\lambda^{\beta}+\delta_{ij}\mathbf{e}^{i}\mathbf{e}^{j}=\delta_{\alpha\beta}\lambda^{\alpha}\lambda^{\beta}+G^{\perp}~{},~{}~{}~{}$ (108)
$H$ $\equiv{1\over 3!}H_{\alpha\beta\gamma}\lambda^{\alpha}\wedge\lambda^{\beta}\wedge\lambda^{\gamma}+{1\over 2}H_{ij\alpha}\mathbf{e}^{i}\wedge\mathbf{e}^{j}\wedge\lambda^{\alpha}+{1\over 3!}H_{ijk}\mathbf{e}^{i}\wedge\mathbf{e}^{j}\wedge\mathbf{e}^{k}$ (109)
$=CS(\lambda)+H^{\perp}~{},$ (110)
where $H_{\alpha\beta\gamma}$ are the structure constants of $\mathfrak{g}$, $CS(\lambda)$ is the Chern-Simons form of $\lambda$, and $G^{\perp}=\delta_{ij}\mathbf{e}^{i}\mathbf{e}^{j}$ and $H^{\perp}={1\over 3!}H_{ijk}\mathbf{e}^{i}\wedge\mathbf{e}^{j}\wedge\mathbf{e}^{k}$ are the components of the metric and torsion orthogonal to all directions spanned by the vector fields generated by $\mathfrak{g}$ on $M^{2n}$. The component $H_{\alpha\beta i}$ of the torsion vanishes because of the closure of the algebra $\mathfrak{g}$ of $\widehat{\nabla}$-covariantly constant vector fields. Note that the Chern-Simon form of $\lambda$ is defined as
$CS(\lambda)\equiv{1\over 3}\delta_{\alpha\beta}\lambda^{\alpha}\wedge d\lambda^{\beta}+{2\over 3}\delta_{\alpha\beta}\lambda^{\alpha}\wedge\mathcal{F}^{\beta}~{},$ (111)
where
${\mathcal{F}}^{\alpha}\equiv d\lambda^{\alpha}-{1\over 2}H^{\alpha}{}_{\beta\gamma}\lambda^{\beta}\wedge\lambda^{\gamma}={1\over 2}H^{\alpha}{}_{ij}\mathbf{e}^{i}\wedge\mathbf{e}^{j}~{},$ (112)
is the curvature of $\lambda$. Using these definitions, the last line of (110) can be verified by a direct calculation. A consequence of (110) is
${\mathrm{d}}H={\mathrm{d}}H^{\perp}+\delta_{\alpha\beta}\mathcal{F}^{\alpha}\wedge\mathcal{F}^{\beta}~{},$ (113)
which expresses the exterior derivative of $H$ in terms of that $H^{\perp}$ and $\mathcal{F}$.
Before we proceed to describe the Hermitian structure, the components of the frame connection $\widehat{\Omega}$ of $\widehat{\nabla}$ in the frame $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$ satisfy the conditions
$\widehat{\Omega}_{A}{}^{\alpha}{}_{B}=\Omega_{A}{}^{\alpha}{}_{B}+\frac{1}{2}H^{\alpha}{}_{AB}=0~{},$ (114)
where $\Omega$ are the components of the frame connection of the Levi-Civita connection $D$, and the index $A=(\alpha,i)$ and similarly for $B$. This can be solved to express some of the components of $\Omega$ in terms of $H$ as
$\Omega_{\alpha\beta\gamma}=\frac{1}{2}H_{\alpha\beta\gamma}~{},~{}~{}~{}\Omega_{i\alpha j}=-\frac{1}{2}H_{\alpha ij}~{},~{}~{}~{}\Omega_{\alpha i\beta}=\Omega_{i\alpha\beta}=0~{}.$ (115)
Some further simplification is possible upon a more careful choice of the frame $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$. For this notice that $\iota_{\widehat{V}_{\alpha}}G^{\perp}=0$ and ${\mathcal{L}}_{\widehat{V}_{\alpha}}G^{\perp}=0$. Thus, $G^{\perp}$ is orthogonal to the directions spanned by the vector fields $\widehat{V}_{\alpha}$ and $\widehat{V}^{K}\partial_{K}G^{\perp}_{IJ}=0$. As a result, one can always choose a frame $\mathbf{e}^{i}$ such that
$V^{J}_{a}\partial_{J}\mathbf{e}_{I}^{i}=0~{}.$ (116)
Then, the torsion free condition for the Levi-Civita frame connection implies that
$\Omega_{\alpha ij}=-\Omega_{ij\alpha}=\frac{1}{2}H_{\alpha ij}~{}.$ (117)
In some examples that will be presented below, the condition on the frame (116) arises naturally in the construction. In any case, the additional conditions (117) are useful to simplify calculations.
One such computation is the decomposition of the curvature $\widehat{R}$ of $\widehat{\nabla}$ connection in the frame $(\lambda^{\alpha}.\mathbf{e}^{i})$. Using (116) as well as ${\mathcal{L}}_{V_{a}}H=0$ and $\iota_{V_{a}}dH=0$, which follows from (113), one can find that the curvature of $\widehat{\nabla}$ on $M^{D}$ decomposes as
$\widehat{R}_{ijkm}=\widehat{R}^{\perp}{}_{ijkm}-\delta_{\alpha\beta}\,{\mathcal{F}}^{\alpha}_{ij}\,{\mathcal{F}}^{\beta}_{km}~{},~{}~{}~{}\widehat{R}_{\alpha jkm}=\delta_{\alpha\beta}\widehat{\nabla}_{j}^{\perp}{\mathcal{F}}^{\beta}_{km}~{},$ (118)
$\widehat{R}_{\alpha\beta km}=H_{\alpha\beta\gamma}{\mathcal{F}}^{\gamma}_{km}-{\mathcal{F}}_{\alpha ik}{\mathcal{F}}_{\beta j}{}^{k}+{\mathcal{F}}_{\beta ik}{\mathcal{F}}_{\alpha j}{}^{k}~{},$ (119)
where
^$\widehat{\nabla}_{j}^{\perp}{\mathcal{F}}^{\beta}_{km}=\partial_{j}{\mathcal{F}}^{\beta}_{km}-\widehat{\Omega}_{j}{}^{n}{}_{k}{\mathcal{F}}^{\beta}_{nm}--\widehat{\Omega}_{j}{}^{n}{}_{m}{\mathcal{F}}^{\beta}_{kn}~{},$ (120)
and $\widehat{\Omega}_{i}{}^{j}{}_{k}$ and $\widehat{R}^{\perp}$ is the frame connection and the curvature of the connection $\widehat{\nabla}^{\perp}$ that can be constructed from $G^{\perp}$ and $H^{\perp}$, respectively. In the examples described below, $\widehat{\Omega}_{i}{}^{j}{}_{k}$ and $\widehat{R}^{\perp}$ will be the frame connection and curvature of the connection with skew-symmetric torsion of the base space $N^{2n-2p}$.
The decomposition of the geometry in (110) and the subsequent analysis presented applies to all manifolds $M^{D}$ that admit a connection with skew-symmetric torsion $\widehat{\nabla}$ and $\widehat{\nabla}$-covariantly constant vector fields. Turning to KT manifolds $M^{2n}$ a consequence of (107) is that the Hermitian form $I$ can be decomposed as
$I\equiv{1\over 2}I_{\alpha\beta}\lambda^{\alpha}\wedge\lambda^{\beta}+{1\over 2}I_{ij}\mathbf{e}^{i}\wedge\mathbf{e}^{j}={1\over 2}I_{\alpha\beta}\lambda^{\alpha}\wedge\lambda^{\beta}+I^{\perp}~{},$ (121)
where $I_{\alpha\beta}$ are constants because of ^$\widehat{\nabla}I=0$ and $I^{\perp}$ is again orthogonal to all directions spanned by the vector fields generated by $\mathfrak{g}$ on $M^{2n}$. As it has already been mentioned the integrability of the complex structure $I$ implies that $H$ is a $(2,1)\oplus(1,2)$ form on $M^{D}$. In terms of this frame this condition decomposes as
$H_{\delta\alpha\beta}I^{\delta}{}_{\gamma}+H_{\delta\gamma\alpha}I^{\delta}{}_{\beta}+H_{\delta\beta\gamma}I^{\delta}{}_{\alpha}-H_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}I^{\alpha^{\prime}}{}_{\alpha}I^{\beta^{\prime}}{}_{\beta}I^{\gamma^{\prime}}{}_{\gamma}=0~{},$ (122)
${\mathcal{F}}^{\alpha}_{ki}I^{k}{}_{j}-{\mathcal{F}}^{\alpha}_{kj}I^{k}{}_{i}+I^{\alpha}{}_{\beta}({\mathcal{F}}^{\beta}_{mn}I^{m}{}_{i}I^{n}{}_{j}-{\mathcal{F}}^{\beta}_{ij})=0~{},$ (123)
$H_{mij}I^{m}{}_{k}+H_{mki}I^{m}{}_{j}+H_{mjk}I^{m}{}_{i}-H_{i^{\prime}j^{\prime}k^{\prime}}I^{i^{\prime}}{}_{i}I^{j^{\prime}}{}_{j}I^{k^{\prime}}{}_{k}=0~{}.$ (124)
Furthermore, the Lie derivative of the Hermitian form along the vector fields $\widehat{V}_{\alpha}$ in the same frame is
${\mathcal{L}}_{\widehat{V}_{\alpha}}I_{\beta\gamma}=H^{\delta}{}_{\alpha\beta}I_{\delta\gamma}-H^{\delta}{}_{\alpha\gamma}I_{\delta\beta}~{},$ (125)
${\mathcal{L}}_{\widehat{V}_{\alpha}}I_{ij}=-{\mathcal{F}}_{\alpha ki}I^{k}{}_{j}+{\mathcal{F}}_{\alpha kj}I^{k}{}_{i}~{}.$ (126)
It is clear from this that the Hermitian form $I$ is invariant under the action of $\mathfrak{g}$ iff the Hermitian form of the fibre is invariant under the adjoint action $\mathfrak{g}$ and the curvature of $\lambda$ is a (1,1) form. It is well known that the former condition can only be satisfied iff $\mathfrak{g}$ is abelian [^15] One way to see this is to notice that there always exist a compact Lie group $\mathcal{G}$ with Lie algebra $\mathfrak{g}$ as $\mathfrak{g}$ admits a positive definite invariant inner product under the $\mathfrak{g}$-adjoint action. Then, the vanishing of the first condition in (126) implies that the associated Hermitian form on $\mathcal{G}$ is bi-invariant and so covariantly constant with respect to the Levi-Civita connection. Thus $\mathcal{G}$ is Kähler and the only such (compact) group manifolds are tori of even dimension.. Furthermore, the vanishing of the second condition in (126), i.e. the restriction for $\mathcal{F}$ to be a (1,1) form, implies the middle condition for the integrability of the complex structure $I$ in (124). Though, the middle condition for the integrability of the complex structure is weaker than the second condition in (126). This is significant as this allows the construction of examples of KT manifolds for which their Hermitian form is not invariant under a group action – the invariance of the Hermitian form under the action of $\mathfrak{g}$ may not be one of the requirements of the construction. It remains to give the Lee form of the KT geometry. This is given as
$\widehat{\theta}$ $\equiv D^{P}I_{PQ}I^{Q}{}_{K}dx^{K}=-\frac{1}{2}I^{PQ}H_{PQL}I^{L}{}_{K}dx^{K}$ (127)
$=-\frac{1}{2}H_{\beta\gamma\delta}I^{\beta\gamma}I^{\delta}{}_{\alpha}\lambda^{\alpha}-\frac{1}{2}I^{\beta}{}_{\alpha}\mathcal{F}_{\beta ij}I^{ij}\lambda^{\alpha}-\frac{1}{2}H_{mnk}I^{mn}I^{k}{}_{i}\mathbf{e}^{i}~{},$ (128)
where the last term can be thought of as the Lee form of $G^{\perp}.H^{\perp}$ and $I^{\perp}$.
To begin let us construct examples of KT manifolds with $\widehat{\nabla}$-covariantly constant vector fields. For this consider principal bundles with fibre a group manifold $\mathcal{G}$ of dimension $2p$ that admits a left invariant KT structure, a principal bundle connection $\lambda$ and base space $N^{2n-2p}$ a KT manifold. In particular, the left invariant KT structure on $\mathcal{G}$ is described by the metric $G^{\mathcal{G}}$, torsion 3-form $H^{\mathcal{G}}$ and Hermitian form $I^{\mathcal{G}}$ as
$G^{\mathcal{G}}=\delta_{\alpha\beta}L^{\alpha}L^{\beta}~{},~{}~{}~{}H^{\mathcal{G}}=\frac{1}{3!}H_{\alpha\beta\gamma}L^{\alpha}\wedge L^{\beta}\wedge L^{\gamma}~{},$ (129)
$I^{\mathcal{G}}=\frac{1}{2}I_{\alpha\beta}L^{\alpha}\wedge L^{\beta}~{},$ (130)
respectively, where $G^{\mathcal{G}}$ and $H^{\mathcal{G}}$ are bi-invariant, i.e. invariant under both the left and right actions of $\mathcal{G}$ on itself, and $(L^{\alpha},\alpha=1,\dots,2p)$ is a left invariant (co-)frame on $\mathcal{G}$. The torsion $H^{\mathcal{G}}$ is given by the structure constants of $\mathcal{G}$. The components $G^{\perp},H^{\perp}$ and $I^{\perp}$ of the metric, torsion and Hermitian form on $M^{2n}$ are pull-back of those on $N^{2n-2p}$. The Hermitian structure $I$ on $M^{2n}$ can be written as in (121) using the components $I_{\alpha\beta}$ of the Hermitian form on the fibre but now taken with respect to the connection $\lambda$ as well as $I^{\perp}$ from the base space. It turns out that with these data $M^{2n}$ is a KT manifold, i.e. $H=-\iota_{I}{\mathrm{d}}I$, iff the curvature of $\lambda$ is a (1,1)-form on $N^{2n-2p}$. Incidentally, this is also the condition, together with the integrability of the complex structure $I^{\perp}$ on $N^{2n-2p}$ and that on the fibre $\mathcal{G}$, for the complex structure $I$ to be integrable on $M^{2n}$. However, the Hermitian form $I$ is not invariant under the action of $\mathcal{G}$, unless $\mathcal{G}$ is abelian, as the vanishing of first condition in (126) required for invariance cannot be satisfied.
By convention, on principal bundles $\mathcal{G}$ acts on each fibre from the right – from now on we shall denote this action with $\mathcal{G}_{R}$. The metric $G$ and torsion $H$ on $M^{2n}$, as described in the previous paragraph, are invariant under the action of $\mathcal{G}_{R}$. However, the complex structure $I$ may not be invariant under the same action unless the Hermitian form of the fibre $\mathcal{G}$ is bi-invariant. As it has been already been explained this can only happen iff $\mathcal{G}$ is abelian. Another possibility arises whether $M^{2n}$ is a product $M^{2n}=\mathcal{G}\times N^{2n-2p}$ and $\lambda$ is the trivial connection – in particular, ${\mathcal{F}}=0$. In such a case, $\lambda$ can be identified with the left invariant 1-forms on $\mathcal{G}$ and the metric, torsion and complex structure can be constructed as in (110) and (121), respectively. Therefore, the KT structure on $M^{2n}$ is the trivial sum of that of the fibre with that of the base space. Of course $G$ and $H$ are invariant under $\mathcal{G}_{R}$ but $I$ is not. However, in this case $I$ is invariant under the left action $\mathcal{G}_{L}$ on $\mathcal{G}$$G$ and $H$ are also invariant under $\mathcal{G}_{L}$.
The holonomy of $\widehat{\nabla}$ of the principal bundle KT geometries constructed above will be $U(n-p)$. For applications to scale and conformal invariant sigma models, the holonomy of $\widehat{\nabla}$ has to be restricted to $SU(n-p)$. One way to achieve this is to take the fibre of the principal bundle to be a KT group manifold, as in the KT $U(n-p)$ case, and further restrict the curvature of the principal bundle as
${\mathcal{F}}^{\alpha}_{km}(I^{\perp})^{k}{}_{i}(I^{\perp})^{m}{}_{j}={\mathcal{F}}^{\alpha}_{ij}~{},~{}~{}~{}{\mathcal{F}}^{\alpha}_{ij}(I^{\perp})^{ij}=0~{},$ (131)
i.e. $\lambda$ has to satisfy the $SU(n-p)$ instanton-like condition or equivalently the Hermitian-Einstein condition, see (57), on $N^{2n-2p}$. Moreover, one has to choose the base space $N^{2n-2p}$ to be a CYT manifold. As it can be seen from (119) for such a manifold, the holonomy of $\widehat{\nabla}$ will be included in $SU(n-p)$. Thus $M^{2n}$ will be a CYT manifold. In particular sigma models with such a target space will be scale invariant up to two loops in sigma model perturbation theory.
For sigma models with target space the CYT manifold $M^{2n}$, whose geometry has been described above, to be conformally invariant, the additional condition ^$\widehat{\nabla}(\widehat{\theta}+2{\mathrm{d}}\Phi)=0$ has to be satisfied. The Lee form of $M^{2n}$ can be written as
$\widehat{\theta}=-\frac{1}{2}H_{\beta\gamma\delta}I^{\beta\gamma}I^{\delta}{}_{\alpha}\lambda^{\alpha}-\frac{1}{2}H_{mnk}I^{mn}I^{k}{}_{i}\mathbf{e}^{i}=\widehat{\theta}^{\mathcal{G}}+\widehat{\theta}^{\perp}~{},$ (132)
where $\widehat{\theta}^{\mathcal{G}}$ is the Lee form of $\mathcal{G}$ taken along the vertical directions and $\widehat{\theta}^{\perp}$ is the Lee form of the base space $N^{2n-2p}$. Therefore, the conformal invariance condition ^$\widehat{\nabla}(\widehat{\theta}+2{\mathrm{d}}\Phi)=0$ implies that $N^{2n-2p}$ should satisfy
^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)=0~{},$ (133)
as $\widehat{\theta}^{\mathcal{G}}$ is $\widehat{\nabla}$-covariantly constant. This is a conformal invariance condition for sigma models with target space $N^{2n-2p}$. For example $N^{2n-2p}$ can be chosen to be conformally balanced. In such a case, the second condition for conformal invariance in (93) will also be satisfied.
More examples can be constructed by again starting with fibre a group manifold with a KT structure and by choosing a connection $\lambda$ to satisfy the instanton-like condition
${\mathcal{F}}^{\alpha}_{km}(I^{\perp})^{k}{}_{i}(I^{\perp})^{m}{}_{j}={\mathcal{F}}^{\alpha}_{ij}~{},~{}~{}~{}{\mathcal{F}}^{\alpha}_{ij}(I^{\perp})^{ij}=2\Lambda^{\alpha}~{},$ (134)
where $\Lambda$ is a constant [^16] Note that the Bianchi identity (54) together with the restriction of the holonomy of $\widehat{\nabla}$ to be included in $SU(n-p)$ and $\iota_{\widehat{V}_{\alpha}}dH=0$ imply that ${\mathcal{F}}^{\alpha}_{ij}(I^{\perp})^{ij}$ is constant. . This is the familiar Hermitian-Einstein condition on ${\mathcal{F}}$ but now with a “cosmological constant”. Furthermore, the base space $N^{2n-2p}$ can be chosen to be a KT manifold with metric $G^{\perp}$, 3-form torsion $H^{\perp}$ and Hermitian structure $I^{\perp}$. The KT structure on $M^{2n}$ can be constructed from that on $\mathcal{G}$ and $N^{2n-2p}$ using the connection $\lambda$ in the way described in (110) and (121). For $M^{2n}$ to admit a CYT, the Ricci form $\widehat{\rho}$ of the $\widehat{\nabla}$ connection vanish. An inspection of (134) reveals that the Ricci form $\widehat{\rho}^{\perp}$ of the connection $\widehat{\nabla}^{\perp}$ of $N^{2n-2p}$ must satisfy
$\widehat{\rho}^{\perp}_{ij}=\Lambda_{\alpha}\mathcal{F}_{ij}^{\alpha}~{},~{}~{}~{}$ (135)
and
$H_{\alpha\beta\gamma}\Lambda^{\gamma}=0~{}.$ (136)
The first condition implies that the first Chern class of $N^{2n-2p}$ becomes trivial upon pulling it back on $M^{2n}$ while the second condition implies that $\Lambda$ must be in the centre of the Lie algebra of $\mathcal{G}$. The construction of KT and CYT manifolds for $\mathcal{G}$ a torus and base space a Kähler manifold has been proposed before in poon .
The Lee form of $M^{2n}$ reads as
$\widehat{\theta}=\widehat{\theta}^{\mathcal{G}}+I_{\alpha\beta}\Lambda^{\beta}\lambda^{\alpha}+\widehat{\theta}^{\perp}~{},$ (137)
where $\widehat{\theta}^{\mathcal{G}}$ and $\widehat{\theta}^{\perp}$ have been defined below (132). Therefore, the conformal invariance condition on $M^{2n}$ reduces to (132) as the remaining components of the Lee form of $M^{2n}$ are $\widehat{\nabla}$-covariantly constant. This allows for more examples as the base space can be chosen to be a conformally balanced KT manifold instead of a conformally balanced CYT manifold, previously required, that is more restrictive.
The above constructions solve all the conditions required for $M^{2n}$ to admit an either KT or CYT structure. However to find examples of CYT manifolds that satisfy the conformal condition, a solution of the condition (12) will also be required [^17] One can find connections for the sigma model gauge sector on $M^{2n}$ whose curvature $F$ satisfies all the necessary conditions. For example, one can pull-back bundles $E$ on $M^{2n}$ from $N^{2n-2p}$ that admit a $SU(n-p)$ instanton-like connection. which can be rewritten using (113) as
${\mathrm{d}}H^{\perp}+\delta_{\alpha\beta}\mathcal{F}^{\alpha}\wedge\mathcal{F}^{\beta}=\frac{\alpha^{\prime}}{4}P_{4}(\widetilde{R},F)~{}.$ (138)
This equation can be seen as a condition on the 3-form $H^{\perp}$ on $N^{2n-2p}$ given $\lambda$, $\widetilde{R}$ and $F$.
It remains to comment on the two cases that arise whenever $p=1$ or $p=n$. In the former case, $\mathfrak{g}=\mathbb{R}\oplus\mathbb{R}$ and it is a special case of the geometries described above. If the action of $\mathbb{R}\oplus\mathbb{R}$ can be integrated to a 2-torus action, $T^{2}$, on $M^{2n}$, then $M^{2n}$ must be a $T^{2}$ bundle and many examples can be constructed. In the latter case, $M^{2n}$ will be a group manifold with a left invariant KT structure – the holonomy of $\widehat{\nabla}$ is trivial..
4.3.2 HKT geometry and conformal invariance
Next suppose that $M^{4k}$ is an HKT manifold that satisfies the conditions for conformal invariance with $\widehat{V}\not=0$ – again $\widehat{V}$ is normalised to have length $1$. As for KT manifolds, the conditions ^$\widehat{\nabla}\widehat{V}=0$ and ^$\widehat{\nabla}I_{r}=0$ imply that ^$\widehat{\nabla}\widehat{Z}_{r}=0$, where $\widehat{Z}_{r}^{J}=-I_{r}{}^{J}{}_{K}\widehat{V}^{K}$. The vector fields $\widehat{V}$ and $\widehat{Z}_{r}$ are orthogonal $G(\widehat{V},\widehat{Z}_{r})=0$, $G(\widehat{Z}_{r},\widehat{Z}_{s})=0$ for $r\not=s$, and $\widehat{Z}_{r}$ have length $1$ as well. Furthermore, it can be shown that $\widehat{Z}_{r}$, as $\widehat{V}$, leave invariant $G$, $H$ and $F$ – the latter up to a gauge transformation. Note that $\iota_{\widehat{Z}_{r}}F=0$ as a consequence of $\iota_{\widehat{V}}F=0$ and that $F$ is an (1,1)-form with respect to each $I_{r}$. The proof that $H$ and $F$ are invariant under the action of $\widehat{V}$ and $\widehat{Z}_{r}$ is similar to that given in (101) and (98), respectively. One consequence of ^$\widehat{\nabla}\widehat{V}=\widehat{\nabla}\widehat{Z}_{r}=0$ is that the holonomy of $\widehat{\nabla}$ reduces to a subgroup of $Sp(k-1)$.
To continue, let $\mathfrak{g}$ be the Lie algebra spanned by $\widehat{V}$ and $\widehat{Z}_{r}$ as well as all their commutators including the vector fields constructed by acting with $I_{r}$ on the commutators. All these vector fields, denoted by $\widehat{V}_{a}$, will be $\widehat{\nabla}$ covariantly constant, $\iota_{\widehat{V}_{a}}F=0$ and leave invariant $G$, $H$ and $F$. To proof of the invariance of $G$, $H$ and $F$ follows from the arguments presented in the KT case and in particular those in (101), (98), (102), (103 and (105). As in the KT case the metric and torsion can be written as in (110) but now $\mathfrak{g}$ has dimension $4q$ and the holonomy of $\widehat{\nabla}$ reduces to $Sp(k-q)$. Furthermore, the Hermitian forms can be written as
$I_{r}=\frac{1}{2}I_{r\alpha\beta}\lambda^{\alpha}\wedge\lambda^{\beta}+\frac{1}{2}I_{rij}\mathbf{e}^{i}\wedge\mathbf{e}^{j}=\frac{1}{2}I_{r\alpha\beta}\lambda^{\alpha}\wedge\lambda^{\beta}+I_{r}^{\perp}~{}.$ (139)
These will not be invariant under the action of $\mathfrak{g}$ unless $\mathfrak{g}$ is abelian. If $q=k$, then $M^{4k}$ is a group manifold with a left invariant HKT structure Spindel ; OP . For $q=1$, there are two possibilities depending on whether $\mathfrak{g}$ is identified with $\oplus^{4}\mathbb{R}$ or $\mathbb{R}\oplus\mathfrak{su}(2)$. In both cases, the metric and torsion are invariant under $\mathcal{G}_{R}$ and they are expressed as in (110).
To construct examples of HKT geometries that solve the conditions for conformal invariance consider principal bundles with fibre a group manifold $\mathcal{G}$ with a left invariant HKT structure equipped with a connection $\lambda$ over a base space $N^{4k-4q}$ whose geometry will be described later. The left invariant HKT structure on $\mathcal{G}$ with metric $G^{\mathcal{G}}$, torsion $H^{\mathcal{G}}$ and Hermitian forms $I_{r}^{\mathcal{G}}$ is
$G^{\mathcal{G}}=\delta_{\alpha\beta}L^{\alpha}L^{\beta}~{},~{}~{}~{}H^{\mathcal{G}}=\frac{1}{3!}H_{\alpha\beta\gamma}L^{\alpha}\wedge L^{\beta}\wedge L^{\gamma}~{},$ (140)
$I_{r}^{\mathcal{G}}=\frac{1}{2}I_{r\alpha\beta}L^{\alpha}\wedge L^{\beta}~{},$ (141)
respectively, where $G^{\mathcal{G}}$ and $H^{\mathcal{G}}$ are bi-invariant, i.e. invariant under both the left and right actions of $\mathcal{G}$ on itself, and $(L^{\alpha},\alpha=1,\dots,2p)$ is a left invariant (co-)frame on $\mathcal{G}$.
One possibility is to choose the base manifold $N^{4(k-q)}$ to admit an HKT structure with metric $G^{\perp}$, torsion $H^{\perp}$ and Hermitian forms $I_{r}^{\perp}$. If in addition the connection $\lambda$ of the principal bundle is chosen such that the curvature $\mathcal{F}$ is a (1,1) form with respect to all three complex structures $I_{r}^{\perp}$ on $N^{4(k-q)}$, i.e.
$\mathcal{F}^{\alpha}_{mn}(I_{r}^{\perp})^{m}{}_{i}(I_{r}^{\perp})^{n}{}_{j}=\mathcal{F}^{\alpha}_{ij}~{},~{}~{}~{}r=1,2,3~{},~{}~{}\mathrm{no~{}summation}$ (142)
then $M^{2n}$ with metric $G$ and torsion $H$ given in (110) and Hermitian forms $I_{r}$ given in (139) will be an HKT manifold. A related construction for HKT manifolds has been suggested before by Verbitsky . The Lee form of such HKT manifold is
$\widehat{\theta}=\widehat{\theta}^{\mathcal{G}}+\widehat{\theta}^{\perp}~{},$ (143)
where $\widehat{\theta}^{\mathcal{G}}$ is the Lee form of ${\mathcal{G}}$ evaluated along the vertical directions, i.e. $\widehat{\theta}^{\mathcal{G}}=\widehat{\theta}^{\mathcal{G}}_{\alpha}\lambda^{\alpha}$ instead of $\widehat{\theta}^{\mathcal{G}}=\widehat{\theta}^{\mathcal{G}}_{\alpha}L^{\alpha}$, and $\widehat{\theta}^{\perp}$ is the Lee form of $N^{4k-4q}$. As $\widehat{\theta}^{\mathcal{G}}$ is $\widehat{\nabla}$-covariantly constant, the condition for conformal invariance becomes
^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)=0~{},$ (144)
i.e. it reduces to demanding that the base space $N^{4k-4q}$ satisfies the condition for conformal invariance. This can be solved by taking $N^{4k-4q}$ to be conformally balanced.
To construct another class of geometries that can satisfy the conformal invariance condition let us demand that the Hermitian forms transform under $\mathfrak{g}$ as
$\mathcal{L}_{\widehat{V}_{\alpha}}I_{r}=(B_{\alpha})^{s}{}_{r}I_{s}~{},$ (145)
where $B_{\alpha}$ are constant $3\times 3$ matrices. These mtrices are restricted as follows. Taking the Lie derivative of the relation hyper-complex structure relation $I_{r}I_{s}=-\delta_{rs}{\mathbf{1}}+\epsilon_{rs}{}^{t}I_{t}$, one concludes
$(B_{\alpha})_{rs}=-(B_{\alpha})_{sr}~{},~{}~{}~{}(B_{\alpha})_{rs}=\delta_{rt}(B_{\alpha})^{t}{}_{s}~{}.$ (146)
Thus one can write $(B_{\alpha})_{rs}=B^{t}_{\alpha}\epsilon_{trs}$. Furthermore, taking the Lie derivative of (145) with respect to $\mathcal{L}_{\widehat{V}_{\beta}}$ and using $[\mathcal{L}_{\widehat{V}_{\alpha}},\mathcal{L}_{\widehat{V}_{\beta}}]=\mathcal{L}_{[\widehat{V}_{\alpha},\widehat{V}_{\beta}]}$, we find that
$[B_{\alpha},B_{\beta}]^{r}{}_{s}=-H_{\alpha\beta}{}^{\gamma}(B_{\gamma})^{r}{}_{s}~{},$ (147)
and so the matrices $B_{\alpha}$ are a 3-dimensional representation of $\mathfrak{g}$. Next suppose that we again consider a principal bundle $M^{4k}$ with fibre an HKT group manifold $\mathcal{G}$ with Lie algebra $\mathfrak{g}$ equipped with a connection $\lambda$ and base space $N^{4k-4q}$ whose geometry will be determined below. Using the frame $(\lambda^{\alpha},\mathbf{e}^{i})$ on $M^{4k}$ to decompose the Hermitian forms as (139), the first two integrability conditions in (124) for each the complex structure on $M^{4k}$ yield
$I_{r}{}^{\beta}{}_{\alpha}(A_{\beta})^{s}{}_{r}+\epsilon_{rt}{}^{s}(B_{\alpha})^{t}{}_{r}=0~{},$ (148)
where the index $r$ should not be summed over. These are additional conditions on the representation $(B_{\alpha})$ of $\mathfrak{g}$. The last condition in (124) is independent from the remaining two and has to be satisfied as well. It imposes the integrability of $I_{r}$ on the base space $N^{4k-4q}$. It remains to investigate the properties of the connection $\lambda$ and to determine the geometry of the base space. For the former, notice from (145) that
$\mathcal{L}_{\widehat{V}_{\alpha}}(I_{r})_{ij}\equiv-\mathcal{F}_{\alpha ki}{}I_{r}{}^{k}{}_{j}+\mathcal{F}_{\alpha kj}{}I_{r}{}^{k}{}_{i}=(B_{\alpha})^{s}{}_{r}(I_{s})_{ij}~{}.$ (149)
This implies that $\lambda$ is a $\mathfrak{sp}(k)\oplus\mathfrak{sp}(1)$ instanton. The curvature decomposes as
$\mathcal{F}^{\alpha}=\mathcal{F}_{\mathfrak{sp}(k)}^{\alpha}+\mathcal{F}_{\mathfrak{sp}(1)}^{\alpha}~{},$ (150)
where the $\mathcal{F}_{\mathfrak{sp}(k)}$ component is not restricted by (149) as it is projected out while the $\mathcal{F}_{\mathfrak{sp}(1)}$ component is determined in terms of representation $(B_{\alpha})$ and the Hermitian forms of the base space $I_{r}^{\perp}$. One way to determine the geometry of the base space is to find the restrictions that the Hermitian forms $I^{\perp}$ satisfy on $N^{4k-4q}$. For this consider a local section of the principal bundle and pull back the frame covariant derivative with torsion of $I^{\perp}$ onto $N^{4k-4q}$. If $(y^{\mu},\mu=1,\dots,4k-4q)$ are some coordinates on $N^{4k-4q}$, one has that
^$\widehat{\nabla}_{\mu}I_{rij}$ $=\partial_{\mu}I_{rij}-\mathbf{e}_{\mu}^{m}\widehat{\Omega}_{m}{}^{k}{}_{i}I_{rkj}+\mathbf{e}_{\mu}^{m}\widehat{\Omega}_{m}{}^{k}{}_{j}I_{rki}-\lambda_{\mu}^{\alpha}\widehat{\Omega}_{\alpha}{}^{k}{}_{i}I_{rkj}+\lambda_{\mu}^{\alpha}\widehat{\Omega}_{\alpha}{}^{k}{}_{j}I_{rki}$ (151)
=^$=\widehat{\nabla}_{\mu}^{\perp}I_{rij}+\lambda_{\mu}^{\alpha}\mathcal{F}_{\alpha}{}^{k}{}_{i}I_{rkj}-\lambda_{\mu}^{\alpha}\mathcal{F}_{\alpha}{}^{k}{}_{j}I_{rki}$ (152)
=^$=\widehat{\nabla}_{\mu}^{\perp}I_{rij}-\lambda_{\mu}^{\alpha}(B_{\alpha})^{s}{}_{r}I_{sij}=0~{}.$ (153)
where ^$\widehat{\nabla}_{\mu}^{\perp}$ is the frame connection of $N^{4k-4q}$ with torsion $H^{\perp}$. Therefore, $N^{4k-4q}$ admits a Quaternionic-Kähler with torsion (QKT) structure [^18] A QKT manifold admits locally three complex structures $I_{r}$ that satisfy the algebra of imaginary unit quaternions and the covariant constancy condition ^$\widehat{\nabla}I_{r}-Q^{s}{}_{r}I_{t}=0$, where $\widehat{\nabla}$ is the metric connection with torsion a 3-form $H$ and $Q$ is an $\mathfrak{sp}(1)$ connection. which has been investigated before in HPO . Reversing the construction one can start from a QKT manifold $N^{4k-4q}$ with torsion a $(2,1)\oplus(1,2)$ form with respect to all complex structures $I_{r}^{\perp}$ such that
^$\widehat{\nabla}_{\mu}^{\perp}I_{rij}-Q_{\mu}{}^{s}{}_{r}I_{sij}=0~{}.$ (154)
Then choose an HKT group manifold $\mathcal{G}$ that admits a 3-dimensional orthogonal representation $(B_{\alpha})$ that satisfies the conditions (148). Next construct a principal bundle $M^{4k}$ with fibre $\mathcal{G}$ that admits a connection $\lambda$ that is a $\mathfrak{sp}(k)\oplus\mathfrak{sp}(1)$ instanton and set $\lambda_{\mu}^{\alpha}(B_{\alpha})^{s}{}_{r}=Q_{\mu}{}^{s}{}_{r}$. This will solve all the conditions for $M^{4k}$ to admit a HKT structure with $\widehat{\nabla}$-covariantly constant vector fields generated by the $\mathcal{G}$ group action. The only remaining condition is to impose (138) on $H^{\perp}$. To investigate whether such manifolds satisfy the condition for conformal invariant, one has to compute the Lee form. It turns out that such manifolds satisfy the conformal condition provided that (144) holds. This can be satisfied for conformally balanced QKT manifold $N^{4k-4q}$.
4.3.3 $G_{2}$ and $\mathrm{Spin}(7)$ geometry and conformal invariance
If the connection $\widehat{\nabla}$ on $M^{7}$ has $G_{2}$ holonomy, the existence of an additional $\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ will reduce the holonomy to a subgroup of $SU(3)$. This is because the isotropy group of a vector in the 7-dimensional representation of $G_{2}$ is $SU(3)$. The $SU(3)$ structure is associated with the (almost) “Hermitian” form $\iota_{\widehat{V}}\varphi$ and the $(3,0)\oplus(0,3)$ form $\iota_{\widehat{V}}{}^{*}\varphi$, where $\varphi$ is the $\widehat{\nabla}$-covariantly constant form $G_{2}$ on $M^{7}$. Of course both $\iota_{\widehat{V}}\varphi$ and $\iota_{\widehat{V}}{}^{*}\varphi$ are $\widehat{\nabla}$-covariantly constant. Examples can be constructed as circle principal fibrations over CYT 6-dimensional manifolds $N^{6}$ with a principal bundle connection $\lambda$ an $SU(3)$ instanton. The metric and torsion on $M^{7}$ can be constructed as in (110), where $\alpha$ takes value $\alpha=0$. Moreover, the $G_{2}$ form on $M^{7}$ is
$\varphi=\lambda^{0}\wedge I^{\perp}+\chi^{\perp}~{},$ (155)
where $I^{\perp}$ and $\chi^{\perp}$ are the Hermitian and $(3,0)\oplus(0,3)$ of the CYT manifold $M^{6}$ pulled back on $M^{7}$. One can demonstrate that $\mathcal{L}_{\widehat{V}^{0}}\varphi=0$. The Lee form $\widehat{\theta}=\widehat{\theta}^{\perp}$, where $\widehat{\theta}^{\perp}$ is the Lee form of $N^{6}$. Thus the condition for conformal invariance becomes ^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2d\Phi)=0$, Therefore the condition for conformal invariance can also be satisfied for $N^{6}$ a conformally balanced CYT manifold.
Similarly, if the connection $\widehat{\nabla}$ on $M^{8}$ has $\mathrm{Spin}(7)$ holonomy, the existence of an additional $\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ will reduce the holonomy to a subgroup of $G_{2}$. This is because the isotropy group of a vector in the 8-dimensional spinor representation of $\mathrm{Spin}(7)$ is $G_{2}$. The $G_{2}$ structure on $M^{8}$ is associated with the 3-form $\iota_{\widehat{V}}\phi$, where $\phi$ is the $\widehat{\nabla}$-covariantly constant self-dual form on $M^{8}$ associated with the $\mathrm{Spin}(7)$ structure. Clearly, $\iota_{\widehat{V}}\phi$ is also $\widehat{\nabla}$-covariantly constant. as in the $G_{2}$ case above, examples can be constructed as circle principal fibrations over conformally balanced $G_{2}$ 7-dimensional manifolds $N^{7}$ with torsion equipped with a connection $\lambda^{0}$, $(\alpha=0)$, a $G_{2}$ instanton. The metric $G$ and torsion $H$ on $M^{8}$ can be constructed as in (110). The self-dual 4-form on $M^{8}$ can be written as
$\phi=\lambda^{0}\wedge\varphi^{\perp}+\star_{7}\varphi^{\perp}~{},$ (156)
where $\varphi^{\perp}$ is the fundamental $G_{2}$ form on $N^{7}$ and $\star_{7}\varphi^{\perp}$ is its dual pulled back on $M^{8}$. Note that $\phi$ is invariant under the action of $\widehat{V}^{0}$, $\mathcal{L}_{\widehat{V}^{0}}\phi=0$. The Lee form $\widehat{\theta}$ of $M^{8}$ is given in terms of $\widehat{\theta}^{\perp}$ on $N^{7}$ associated with the $G_{2}$ of the latter, i.e. $\widehat{\theta}=\widehat{\theta}^{\perp}$. Thus the conformal invariance condition becomes ^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2d\Phi)=0$ and so it is satisfied if $N^{7}$ is conformally balanced as previously stated. Of course in order to construct solutions, in both $G_{2}$ and $\mathrm{Spin}(7)$ cases, the condition (12), which can again be rewritten as (138) with $\lambda$ abelian, has also to be satisfied. More general examples of $G_{2}$ and $\mathrm{Spin}(7)$ geometries that satisfy the conformal invariance conditions can be constructed and these will be reported elsewhere.
Acknowledgements: I thank Edward Witten for many insightful discussions.